### v1

# Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Long Chen1†

Abstract
Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning, but remain limited in visual relation understanding (e.g., scene graph generation), particularly in modeling N-ary relationships that identify multiple semantic roles among an action event. Such a lack of semantic dependencies modeling among multi-entities leads to unreliable outputs, intensifying MLLMs’ hallucinations and over-reliance on language priors. To this end, we propose Relation-R1, the first unified relational comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) within a reinforcement learning (RL) paradigm. Specifically, we first establish foundational reasoning capabilities via SFT, enforcing structured outputs with thinking processes. Then, GRPO is utilized to refine these outputs via multi-reward optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability. Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
11footnotetext: * The first two authors contribute equally to this work.22footnotetext: † Corresponding Author: Long Chen.
1 Introduction
Recent advancements in MLLMs have significantly enhanced holistic image understanding and object-level grounding capabilities, enabling tasks like region captioning and referring question answering. However, current MLLMs exhibit critical limitations in relational scene understanding — a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
Figure 1: Illustrations of our Relation-R1 and other MLLMs. (a) Grounded MLLMs: Link objects in text to image regions but fail to capture object relations. (b) Relation Comprehension MLLMs: Model pairwise object relations but lack generalization to N-ary interactions. (c) Unified Relation Comprehension MLLMs: Jointly handle binary and $N$-ary relation detection.
Despite existing studies have made progress in binary relation detection [1, 2], i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), they fail to capture $N$-nary relationship, that identify $N$-th entities engaged in the relation with their roles (what, who, and where concepts). It leads to neglecting the semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, which may result in unreasonable relation triplets, e.g., “child-drinking-glass”. Furthermore, these unreasonable pair-wise relations may make the MLLMs suffer from both hallucinations and over-relying on language priors rather than visual-semantic evidence. Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings. This raises a critical question: How can we devise a unified framework for joint binary and N-ary relational reasoning with grounded cues?
Intuitively, such complicated relational comprehension requires both reasoning and generalization abilities. As for reasoning, the model must perform multi-step and multi-modal reasoning, it includes entity identification, relation recognition, role inference (e.g., “glass” as a container), localization, and contextual integration. In terms of generalization, models trained on limited data struggle to extrapolate to novel entity combinations or ambiguous contexts (e.g., “cup” as a container). In addition, language priors (e.g., common-sense assumptions) often dominate over visual grounding, further limiting their robustness.
Fortunately, DeepSeek-R1-Zero has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL. Inspired by this breakthrough, we propose a unified relation comprehension framework Relation-R1, that integrates Supervised Fine-Tuning (SFT) and RL to empower MLLMs with relational reasoning and generalization capability. While SFT alone suffers from poor generalization due to reliance on fixed training patterns, RL struggles to produce structured outputs with consistent formatting [3]. To alleviate these limitations, we design a two-stage pipeline:
Stage 1. SFT: Establishes foundational reasoning through cognitive chain-of-thought guidance. The model generates structured outputs by analyzing visual evidence (e.g., object detection, spatial localization, etc.), ensuring visual grounding over language priors.
Stage 2. RL: Refines outputs via GRPO guided by multi-rewards: (1) format rewards, (2) binary relation rewards, and (3) N-ary relation rewards. This stage employs policy gradient updates to prioritize visual-semantic coherence over language priors, fostering robust relational reasoning and generalization.
By first leveraging SFT to establish reasoning foundations and then applying GRPO to optimize for structured, grounded outputs, Relation-R1 achieves robust relational reasoning and generalization to novel contexts, directly addressing the limitations of prior MLLMs.
We validate the effectiveness of Relation-R1 on both the binary relation detection task (e.g., Scene Graph Generation) and N-ary relation detection task (e.g., Grounded Situation Recognition) under both few-shot and full-data settings. Experiments on PSG [4] dataset and SWiG [5] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts. In summary, our contributions are threefold:
  • We incorporate both binary relation detection and N-ary relation detection in the unified relation comprehension framework.
  • We propose Relation-R1, a unified framework that integrates cognitive CoT guided SFT with RL. To the best of our knowledge, this is the first R1-like work to enable both relational reasoning and robust generalization.
  • Experiments on the prevalent relation understanding datasets under different settings demonstrate the effectiveness of Relation-R1.
2 Related Work
Reasoning in Multimodal Large Language Models. Recent Large Language Models (LLMs) have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [6, 7]. A pivotal breakthrough, DeepSeek-R1 [7], leverages large-scale reinforcement learning (RL) with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains. Multimodal Large Language Models (MLLMs) extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [8, 9, 10] For multimodal mathematical reasoning, frameworks such as R1-Onevision [8], Virgo [9], and MM-Eureka [11] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [10] combines cold-start initialization with GRPO to refine complex CoT reasoning. For fine-grained classification and grounding tasks, Visual-RFT [12] and R1-Omni [13] utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning. As for the pixel-level understanding task, SegZero [14] employs reinforcement learning to achieve high-resolution semantic segmentation. Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts. To bridge this gap, our Relation-R1 explicitly integrates relational reasoning into MLLMs via reinforcement learning, enabling systematic inference for both scene interpretation and event prediction.
Scene Graph Generation (SGG). SGG is a critical task in scene understanding, with prior approaches falling into two categories: 1) Two-stage SGG, which sequentially detects objects and infers pairwise relations but suffers from error propagation [15, 16, 17, 18]; 2) One-stage SGG, which unifies detection and relation prediction in end-to-end frameworks (e.g., DETR-based methods [19, 20, 21]) but often lacks multi-role involved relation modeling. Recent efforts in open-vocabulary SGG (OVSGG) leverage Vision-Language Models (VLMs) [22, 23, 24] or MLLMs [1, 2, 25] to handle novel entities/relations, yet limited modeling of complex multi-entity interactions, and suffer from overfitting in base categories during SFT. Our Relation-R1 addresses OVSGG challenges by integrating a cognitive CoT-guided RL framework to model complex multi-entity interactions and mitigate SFT overfitting, enabling robust zero-shot reasoning without predefined category constraints.
Grounded Situation Recognition (GSR). GSR extends scene understanding by jointly modeling actions (verbs) and their associated semantic roles. Pratt et al. [5] pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization. Subsequent works [26, 27, 28, 29] improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations. Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution [15, 16, 17], potentially limiting real-world applicability. While LEX [30] realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions. This paper proposes Relation-R, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
3 Methodology
Figure 2: Overview of the proposed Relation-R1 framework. (a) SFT with cognitive CoT, which enforces format alignment through task-specific reasoning processes. (b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-nary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
3.1 Preliminaries
Binary Relation Detection. This task requires the model to generate an open-ended scene graph given an image, explicitly marking semantic relationships between pairwise objects [1]. The output is structured as a sequence of relation triplets annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates < box> to localize the entity in the image, e.g., <ref>person</red></box>[[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]]</box>. Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes.
N-ary Relation Detection. Similarly, this task aims to output a open-ended grounded situation frame [5] structured as a primary activity (relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates, each specifying an entity’s role, e.g., <agent>person</agent></box>[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]</box>. Each role explicitly defines the entity’s participation in the activity.
3.2 Supervised Fine-Tuning
As mentioned above, directly adopting GRPO faces challenges in generating format-inconsistent scene graph captions and grounded situation frame outputs. To this end, we adopt supervised fine-tuning that serves as a foundational stage for task-specific format alignment (c.f., Figure 2(a)).
However, utilizing merely the answer format as the SFT target risks compromising the model’s reasoning capability by encouraging rigid adherence to syntactic structures [3]. To mitigate this, we devise the cognitive Chain-of-Thoughts (CoT), which incorporates stepwise cognitive processes such as object classification, object grounding, visual relationship inference, etc. This ensures the model retains its ability to perform multi-step reasoning while adhering to output formats.
To avoid the MLLM overfitting to specific CoT patterns during SFT, we devise a cognitive CoT generation prompt that explicitly guides an extra strong MLLM (e.g., Qwen 2.5-VL [31]) to generate step-by-step reasoning processes. To be specific, this cognitive CoT generation prompt consists of task definition, scene graph caption/grounded situation frame examples, and GT scene graph caption/grounded situation frame of the current image. The former two make the MLLM generate the correct task format. Meanwhile, the latter one anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors. The detailed Cognitive CoT generation prompt is left in the Appendix.
These cognitive CoT are explicitly encapsulated within <think> tags, during SFT. By leveraging this structured training signal, the MLLM learns to generate format-compliant outputs and grounded reasoning processes as an initial capability, which is critical for subsequent CoT-guided generalization under the GRPO.
3.3 Group Relative Policy Optimization
Following the Deepseek-R1 [7], the Group Relative Policy Optimization (GRPO) is adopted as the reinforcement learning (RL) algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b). Distinct from critic-dependent methods [32], GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity [7]. Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{i}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge. The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging following objectives:
$J_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}$ (1)
$\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$
where $\rho_{i}$ = $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ quantifies policy change, $\epsilon$ controls clipping thresholds, and $\beta$ penalizes deviations via the KL divergence term. The advantage score $A_{i}$ is standardized as:
$A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
here, $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
$D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1,$ (3)
which ensures controlled exploration without excessive divergence from $\pi_{\theta_{\text{ref}}}$.
The reward function $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding. It incorporates rule-based rewards, which consist of three crucial components: Format Compliance Rewards, Binary Relation Rewards, and $N$-nary Relation Rewards. The former enforces syntactic validity by penalizing outputs that deviate from predefined structural constraints. The latter two components focus on task-specific performance.
Format Rewards. This reward enforces strict adherence to the reasoning template format: <think> thinking process </think><answer> answer </answer>. The reward function is defined as:
$r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
This binary mechanism ensures syntactic compliance while allowing content flexibility.
Binary Relation Rewards. Inspired by established evaluation practices [4, 1], we devise the sample-level triplet Recall and sample-level mean Recall (mRecall) for each predicate category in the response of each sample. Specifically, a triplet is deemed correct if:
  • Triplet Accuracy: All entity categories (subject, predicate, object) in this sample are correctly identified.
  • Spatial Localization: The predicted bounding boxes for the subject and object achieve an Intersection over Union (IoU) >= 0.5 with their respective ground-truth boxes.
The final binary relation reward is computed as a weighted combination of these two metrics:
$r_{\text{binary}}(o)=\alpha\cdot Recall+(1-\alpha)\cdot mRecall,$ (5)
where $\alpha$ balances the trade-off between two metrics.
N-ary Relation Rewards. Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary Relation Rewards, which assess the model’s ability to detect primary activity and multi-entity interactions. Specifically, we calculate the verb accuracy $Acc_{v}$ and noun accuracy $Acc_{n}$:
  • Verb Accuracy: The predicted primary activity (verb) in the grounded situation frame is the same as the ground-truth activity.
  • Noun Accuracy: All entities involved in the event must satisfy IoU >=0.5 with their ground-truth bounding boxes, and their roles must align with the event’s semantic structure.
The eventual N-ary relation reward is computed as a weighted combination of these two accuracies:
$r_{\text{n-ary}}(o)=\beta\cdot Acc_{v}+(1-\beta)\cdot Acc_{n},$ (6)
where hyperparameter $\beta$ balances the trade-off between event type precision and multi-entity structural accuracy. During training, we design a multi-task gate that dynamically selects task-specific rewards based on the presence of <ref> tags in the solutions.
4 Experiment
4.1 Experimental Setup
4.1.1 Datasets
PSG. The Panoptic Scene Graph (PSG) dataset [4] is built upon the COCO dataset [33], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing. It integrates 80 “thing” object categories and 53 “stuff” object categories, alongside 56 relation categories. To enable open-ended scene graph generation, we followed ASMv2 [1], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training.
SWiG. The SWiG dataset [5] extends the imSitu dataset [34] by adding extra bounding box annotations for visible semantic roles, with 69.3% of semantic roles localized. Each image is annotated with a verb and a variable number of semantic roles (1$\sim$6 roles per image), where each verb is associated with three verb frames annotated by independent human annotators. The dataset comprises 25,200 testing images spanning 504 verb categories, 190 semantic role categories, and 9,929 noun entity categories. Unlike traditional output structures that merely predict categories and bounding boxes, we constructed grounded situation frames by filling verb-centric templates (e.g., “The AGENT drinks a LIQUID from a CONTAINER”) with three critical components: role tags (e.g., <agent>, entity categories (e.g., child), and coordinates of corresponding bounding boxes in <box> tags.
4.1.2 Evaluation Metrics
Binary Relation Detection. Following the standard SGG evaluation protocol [15, 1], we reported 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object. 2) mean Recall (mRecall): The average recall across all predicate categories.
N-ary Relation Detection. We adopted the same five evaluation metrics as [5]: 1) verb: the accuracy of verb prediction. 2) value: the noun accuracy for individual semantic role. 3) value-all (val-all): Overall noun prediction accuracy across all semantic roles in an event. 4) grounded-value (grnd): Bounding box accuracy for each semantic role, requiring an IoU >= 0.5 between predicted and ground-truth boxes. 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
4.1.3 Implementation Details
We utilized Qwen2.5-VL-3B [31] as our base model. Relation-R1 was trained on an 8xA100 GPU server. During the SFT phase, we employed a total batch size of 64, with 8 samples processed per training step. For the GRPO component, we used a total batch size of 32. The learning rates were set to 1e-5 for SFT and 1e-7 for GRPO, respectively.
Table 1: Performance(%) comparison with various SGG models on the PSG dataset [4].
            Method Model Size Answer Recall mRecall Mean
IMP [35] ${}_{\text{CVPR'17}}$ Close-ended 16.50 6.50 11.50
MOTIFS [16] ${}_{\text{CVPR'18}}$ 20.00 9.10 14.55
VCTree [36] ${}_{\text{CVPR'19}}$ 20.60 9.70 15.15
GPSNet [37] ${}_{\text{CVPR'20}}$ 17.80 7.00 12.4
PSGFormer [4] ${}_{\text{ECCV'22}}$ 18.60 16.70 17.65
TextPSG [38] ${}_{\text{ICCV'23}}$ Open-ended 4.80 - -
ASMv2 [1] ${}_{\text{ECCV'24}}$ 13B 14.20 10.30 12.23
LLaVA-SpaceSGG [2] ${}_{\text{WACV'25}}$ 13B 15.43 13.23 14.33
Relation-R1 (Ours) 3B 19.59 18.30 18.95
Table 2: Performance(%) comparison with various GSR models on the SWiG [5] dataset.
         Method Answer Verb Value Value-all Gnd Value Gnd Value-all
ISL [5] ${}_{\text{ECCV'20}}$ Close-ended 39.36 30.09 18.62 22.73 7.72
JSL [5] ${}_{\text{ECCV'20}}$ 39.94 31.44 18.87 24.86 9.66
GSRTR [26] ${}_{\text{BMVC'21}}$ 40.63 32.15 19.28 25.49 10.10
CoFormer [28] ${}_{\text{CVPR'22}}$ 44.66 35.98 22.22 29.05 12.21
SituFormer [27] ${}_{\text{AAAI'22}}$ 44.20 35.24 21.86 29.22 13.41
GSRFormer [29] ACM ${}_{\text{ACM MM'22}}$ 46.53 37 .48 23.32 31.53 14.23
OpenSU [39] ${}_{\text{ICCVW'23}}$ Open-ended 50.10 41.20 26.56 34.27 15.70
Relation-R1 (Ours) 53.76 43.55 28.23 36.79 26.84
Table 3: Few-Shot GSR performance comparison. $\dagger$ denotes employing the SFT phrase without CoT.
         Method Setting Verb Value Value-all Gnd Value Gnd Value-all
GSRTR [26] ${}_{\text{BMVC'21}}$ One-Shot 0.35 0.10 0.01 0.08 0.00
CoFormer [28] ${}_{\text{CVPR'22}}$ 0.63 0.23 0.02 0.14 0.01
Relation-R1 (Ours) 27.31 10.21 1.43 9.35 1.22
Relation-R1 (Ours) 31.67 15.37 4.15 14.01 3.41
GSRTR [26] ${}_{\text{BMVC'21}}$ 4-Shot 6.95 4. 32 1.83 2.79 0.65
CoFormer [28] ${}_{\text{CVPR'22}}$ 12.39 8.27 3.93 5.28 1.25
Relation-R1 (Ours) 34.88 17.14 3.94 15.63 3.38
Relation-R1 (Ours) 38.02 21.87 7.39 19.64 5.92
(a) SGG.
(b) GSR.
Figure 3: The statistics of rewards during training.
Figure 4: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG. The orange area denotes the predicate distribution of the training set.
Figure 5: Qualitative results of binary relation detection (scene graph generation).
Figure 6: Qualitative results of N-ary relation detection (grounded situation recognition).
4.2 Quantitative Comparison Result
4.2.1 Binary Relation Detection
Open-ended SGG. We evaluated Relation-R1 on PSG, comparing it against various established and recent SGG models, as detailed in Table 1. Positioned within the open-ended generation category, Relation-R1 achieved SOTA performance with the highest mRecall of 18.30% and overall Mean of 18.95%, substantially surpassing prior open-ended methods like LLaVA-SpaceSGG (13.23% mRecall, 14.33% Mean). It also achieved a competitive Recall of 19.59%. Notably, Relation-R1 attained these leading results using a significantly smaller 3B parameter model compared to the 13B models used by recent competitors like ASMv2 and LLaVA-SpaceSGG. In addition, we visualized predicate-level comparison against ASMv2 and the predicate distribution in Figure 4. As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates. These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
4.2.2 N-ary Relation Detection
Open-ended GSR. We evaluated Relation-R1 for N-ary relation detection on SWiG, comparing it against prior work in Table 2. Within the open-ended setting, Relation-R1 established a new SOTA, achieving top performance across all reported metrics: Verb (53.76%), Value (43.55%), Value-all (28.23%), Gnd Value (36.79%), and Gnd Value-all (26.84%). Notably, Relation-R1 significantly surpassed the previous best open-ended model, OpenSU [39], exhibiting substantial gains, particularly with a remarkable +11.14% absolute improvement on the challenging Gnd Value-all metric. This clearly demonstrated Relation-R1’s superior capacity for comprehending and grounding complex N-ary relationships in visual scenes.
Few-Shot GSR. We also evaluated Relation-R1’s few-shot generalization capability (Table 3). In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Gnd Value-all) compared to CoFormer’s [28] 0.63% and 0.01%, respectively. This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Gnd Value-all), significantly surpassing previous methods. These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
4.3 Qualitative Result
Figure 5 and Figure 6 provided qualitative illustrations of Relation-R1. For Binary Relation Detection (Figure 5), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference. For instance, after identifying the central bed, Relation-R1 pinpoints related objects like the rug and chair and explicitly grounds the inferred spatial relationships (on, beside) between them. This systematic approach demonstrates the successful decomposition of the scene into accurate, grounded object-relation pairs. Similarly, for N-ary Relation Detection (Figure 6), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box. This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
4.4 Diagnostic Experiment
To gain more insights into Relation-R1, we conducted a set of comprehensive studies on SGG and GSR.
Reward Statistics. We visualized the variation curves of rewards in the SGG task and the GSR task, respectively, as displayed in Figure 3(a) and Figure 3(b). It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT. With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend.
5 Conclusion
This paper reveals critical limitations of existing MLLMs in visual relation understanding, particularly concerning N-ary relationships and the detrimental effects of language-induced bias. We propose Relation-R1, an advanced, unified framework designed to jointly tackle both binary and N-ary relational comprehension. Our core innovation lies in a two-stage approach: first establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, and subsequently refining the model’s outputs using RL with GRPO. This RL stage, guided by carefully designed multi-rewards, explicitly promotes visual-semantic grounding and robust generalization. Extensive experiments conducted on the challenging PSG and SWiG datasets validate the effectiveness of Relation-R1, demonstrating SOTA performance across both binary and N-ary relation detection tasks. By successfully integrating CoT-guided SFT with R1-inspired RL refinement, Relation-R1 represents a solid step towards equipping MLLMs with more human-like relational reasoning capabilities.

### v2

@@
-Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Long Chen1†
+Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Kwang-Ting Cheng1,2  Long Chen1†
 Abstract
-Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning, but remain limited in visual relation understanding (e.g., scene graph generation), particularly in modeling N-ary relationships that identify multiple semantic roles among an action event.
-Such a lack of semantic dependencies modeling among multi-entities leads to unreliable outputs, intensifying MLLMs’ hallucinations and over-reliance on language priors.
-To this end, we propose Relation-R1, the first unified relational comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) within a reinforcement learning (RL) paradigm.
+Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning.
+However, they remain limited in visual relation understanding, struggling even with binary relation detection, let alone N-ary relations involving multiple semantic roles.
+The core reason is the lack of modeling for structural semantic dependencies among multi-entities, leading to unreliable outputs, hallucinations, and over-reliance on language priors (e.g., defaulting to “person drinks a milk” if a person is merely holding it).
+To this end, we propose Relation-R1, the first unified relation comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided supervised fine-tuning (SFT) and group relative policy optimization (GRPO) within a reinforcement learning (RL) paradigm.
 Specifically, we first establish foundational reasoning capabilities via SFT, enforcing structured outputs with thinking processes.
-Then, GRPO is utilized to refine these outputs via multi-reward optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability.
+Then, GRPO is utilized to refine these outputs via multi-rewards optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability.
+Furthermore, we investigate the impact of various CoT strategies within this framework, demonstrating that a specific-to-general progressive approach in CoT guidance further improves generalization, especially in capturing synonymous N-ary relations.
 Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
-11footnotetext: * The first two authors contribute equally to this work.22footnotetext: † Corresponding Author: Long Chen.
 1 Introduction
-Recent advancements in MLLMs have significantly enhanced holistic image understanding and object-level grounding capabilities, enabling tasks like region captioning and referring question answering.
+Recent advancements in MLLMs have significantly enhanced holistic image understanding [1, 2] and object-level grounding [3, 4, 5, 6, 7] capabilities, enabling tasks like region captioning and referring question answering.
 However, current MLLMs exhibit critical limitations in relational scene understanding — a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
@@
 (c) Unified Relation Comprehension MLLMs: Jointly handle binary and $N$-ary relation detection.
-Despite existing studies have made progress in binary relation detection [1, 2], i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), they fail to capture $N$-nary relationship, that identify $N$-th entities engaged in the relation with their roles (what, who, and where concepts).
-It leads to neglecting the semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, which may result in unreasonable relation triplets, e.g., “child-drinking-glass”.
-Furthermore, these unreasonable pair-wise relations may make the MLLMs suffer from both hallucinations and over-relying on language priors rather than visual-semantic evidence.
+Despite notable progress in binary relation detection, i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), current models [8, 9] struggle to capture N-nary relationships.
+Accurately modeling such relationships requires models to identify multiple N-th entities engaged in the relation with their distinct semantic roles (what, who, and where concepts) [10, 11].
+The neglect of these crucial structural semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, may result in suboptimal relation triplets, e.g., “child-drinking-glass”.
+Furthermore, these unreliable and superficial pair-wise relations may make the MLLMs prone to both hallucinations and over-reliance on language priors rather than visual-semantic cues (e.g., evidence from semantically grounded visual regions).
 Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings.
 This raises a critical question: How can we devise a unified framework for joint binary and N-ary relational reasoning with grounded cues?
-Intuitively, such complicated relational comprehension requires both reasoning and generalization abilities.
-As for reasoning, the model must perform multi-step and multi-modal reasoning, it includes entity identification, relation recognition, role inference (e.g., “glass” as a container), localization, and contextual integration.
-In terms of generalization, models trained on limited data struggle to extrapolate to novel entity combinations or ambiguous contexts (e.g., “cup” as a container).
-In addition, language priors (e.g., common-sense assumptions) often dominate over visual grounding, further limiting their robustness.
-Fortunately, DeepSeek-R1-Zero has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL.
-Inspired by this breakthrough, we propose a unified relation comprehension framework Relation-R1, that integrates Supervised Fine-Tuning (SFT) and RL to empower MLLMs with relational reasoning and generalization capability.
-While SFT alone suffers from poor generalization due to reliance on fixed training patterns, RL struggles to produce structured outputs with consistent formatting [3].
-To alleviate these limitations, we design a two-stage pipeline:
+Intuitively, such complicated relation comprehension requires both strong reasoning and generalization abilities.
+As for reasoning, this capability is essential as complex relations are often implicit and require the model to perform sophisticated inference by integrating diverse and often subtle visual cues.
+This involves a multi-step reasoning process, e.g., entity identification (e.g., recognizing “glass” category appears in the image), relation recognition (e.g., recognizing “drink” action appears in the image), role inference (e.g., “glass” as a container), localization, and contextual integration (e.g., observing a child holding a glass near their mouth suggests the “drink” action is likely occurring).
+In terms of generalization, models trained on limited data struggle to extrapolate to novel interactions and entity combinations or ambiguous contexts (e.g., “cup” as a container for “drinking” action).
+In addition, inductive bias from language priors (i.e., common-sense assumptions learned purely from text that may not always align with the visual reality) often dominates over visual grounding, further limiting their robustness.
+For example, a model might struggle to recognize “drinking” from “milk” if its language training strongly associates this action with “glass”.
+Fortunately, DeepSeek-R1-Zero [12] has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL.
+However, due to the inherent difficulty of reference models in understanding structural relationships, the direct application of RL often struggles to produce consistently formatted outputs that are claimed in the input prompt [13].
+Conversely, when solely adopting supervised fine-tuning, the model suffers from poor generalization due to its overfitting to fixed training patterns and the limited diversity of annotated relational data.
+To this end, inspired by the breakthrough of R1-series [12, 14, 15] in enabling reasoning and the ability of SFT to provide structured guidance, we propose a unified two-stage relation comprehension framework, Relation-R1, which combines the advantages of both supervised fine-tuning and reinforcement learning to empower MLLMs with relational reasoning and generalization capability:
 Stage 1.
-SFT: Establishes foundational reasoning through cognitive chain-of-thought guidance.
-The model generates structured outputs by analyzing visual evidence (e.g., object detection, spatial localization, etc.), ensuring visual grounding over language priors.
+SFT: Establishes foundational reasoning through step-by-step cognitive chain-of-thought guidance.
+This CoT enables the model to break down complex relation comprehension into a sequence of smaller, interpretable steps with visual evidence, such as object detection, spatial localization, etc.
+Thereby, the model learns to think based on grounded visual content over language inductive bias.
+Furthermore, its annotated relation-contained instructions ensure that the model generates outputs in a standardized format, facilitating the subsequent computation of rewards during RL.
 Stage 2.
-RL: Refines outputs via GRPO guided by multi-rewards: (1) format rewards, (2) binary relation rewards, and (3) N-ary relation rewards.
-This stage employs policy gradient updates to prioritize visual-semantic coherence over language priors, fostering robust relational reasoning and generalization.
-By first leveraging SFT to establish reasoning foundations and then applying GRPO to optimize for structured, grounded outputs, Relation-R1 achieves robust relational reasoning and generalization to novel contexts, directly addressing the limitations of prior MLLMs.
-We validate the effectiveness of Relation-R1 on both the binary relation detection task (e.g., Scene Graph Generation) and N-ary relation detection task (e.g., Grounded Situation Recognition) under both few-shot and full-data settings.
-Experiments on PSG [4] dataset and SWiG [5] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts.
-In summary, our contributions are threefold:
-  • We incorporate both binary relation detection and N-ary relation detection in the unified relation comprehension framework.
-  • We propose Relation-R1, a unified framework that integrates cognitive CoT guided SFT with RL.
-To the best of our knowledge, this is the first R1-like work to enable both relational reasoning and robust generalization.
-  • Experiments on the prevalent relation understanding datasets under different settings demonstrate the effectiveness of Relation-R1.
+RL: Refines the structured outputs generated by the SFT model through group relative policy optimization.
+The learning process is guided by multi-rewards: 1) format rewards, which incentivize the generation of thinking process [12]; 2) binary relation rewards, which encourage accurate identification of pairwise relationships between entities; and 3) N-ary relation rewards, which promote the comprehensive understanding of more complex, multi-entity activity.
+This stage employs policy gradient updates to cultivate the model to explore various potential solutions and optimizes its output based on defined multi-rewards, fostering robust relational reasoning and generalization.
+Furthermore, to gain deeper insights into varying cognitive CoTs, we investigate the impact of template-based CoT and MLLM-generated CoT strategies on binary and N-ary relation comprehension, separately.
+Notably, by employing a progressive paradigm that initially adopts the template-based CoTs to guide the normative cognitive reasoning process, and then fine-tunes the model with the variable reasoning pathways introduced by a few MLLM-generated CoTs, our Relation-R1 demonstrates better generalization ability.
+Especially in the N-ary relations, the capacity to explore synonymous relational expressions has emerged.
+We validate the effectiveness of Relation-R1 on both the binary relation comprehension task (e.g., scene graph generation [16, 8]) and N-ary relation comprehension task (e.g., grounded situation recognition [10]) across diverse experimental settings.
+Experiments on the PSG [17] dataset and SWiG [10] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts.
+In summary, our contributions are fourfold:
+  • We reveal the limitations of the existing MLLMs in relation understanding, and incorporate both binary and N-ary relation detection into one unified relation comprehension framework.
+  • We propose Relation-R1, a unified framework that integrates cognitive CoT-guided SFT with RL.
+To the best of our knowledge, this is the first work that enables both relational reasoning and robust generalization in a deep-thinking fashion.
+  • We comparatively analyze template-based and MLLM-generated CoTs, demonstrating that progressively guiding the learning process with CoTs in a specific-to-general manner enhances generalization capabilities, particularly in capturing synonymous N-ary relations.
+  • Experimental results on the prevalent relation understanding datasets (e.g., +6.84$\sim$6.90gains on PSG dataset) under different settings demonstrate the effectiveness of Relation-R1.
 2 Related Work
-Reasoning in Multimodal Large Language Models.
-Recent Large Language Models (LLMs) have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [6, 7].
-A pivotal breakthrough, DeepSeek-R1 [7], leverages large-scale reinforcement learning (RL) with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains.
-Multimodal Large Language Models (MLLMs) extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [8, 9, 10] For multimodal mathematical reasoning, frameworks such as R1-Onevision [8], Virgo [9], and MM-Eureka [11] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [10] combines cold-start initialization with GRPO to refine complex CoT reasoning.
+Reasoning in MLLMs.
+Recent LLMs have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [18, 12].
+A pivotal breakthrough, DeepSeek-R1 [12], leverages large-scale RL with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains.
+MLLMs extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [14, 19, 20] For multimodal mathematical reasoning, frameworks such as R1-Onevision [14], Virgo [19], and MM-Eureka [21] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [20] combines cold-start initialization with GRPO to refine complex CoT reasoning.
 For fine-grained classification and grounding tasks, Visual-RFT [22] and R1-Omni [15] utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning.
@@
 Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts.
-To bridge this gap, our Relation-R1 explicitly integrates relational reasoning into MLLMs via reinforcement learning, enabling systematic inference for both scene interpretation and event prediction.
+To bridge this gap, Relation-R1 explicitly integrates relational reasoning into MLLMs via RL, enabling systematic inference for both scene interpretation and event prediction.
 Scene Graph Generation (SGG).
@@
 While LEX [38] realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions.
-This paper proposes Relation-R, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
+This paper proposes Relation-R1, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
 3 Methodology
 Figure 2: Overview of the proposed Relation-R1 framework.
-(a) SFT with cognitive CoT, which enforces format alignment through task-specific reasoning processes.
+(a) SFT with progressive guidance of cognitive CoT, which enforces task-specific reasoning processes and format alignment.
 (b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-nary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
@@
 Binary Relation Detection.
-This task requires the model to generate an open-ended scene graph given an image, explicitly marking semantic relationships between pairwise objects [1].
-The output is structured as a sequence of relation triplets annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates < box> to localize the entity in the image, e.g., <ref>person</red></box>[[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]]</box>.
+This task requires the model to generate an open-ended scene graph caption given an image, explicitly marking semantic relationships between pairwise objects [8].
+The output is a triplet-contained caption annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates in < box> tags to localize the entity in the image, e.g., <ref>person</ref><box>[[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]]</box>.
 Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes.
 N-ary Relation Detection.
-Similarly, this task aims to output a open-ended grounded situation frame [5] structured as a primary activity (relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates, each specifying an entity’s role, e.g., <agent>person</agent></box>[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]</box>.
-Each role explicitly defines the entity’s participation in the activity.
+Similarly, this task aims to output a open-ended grounded situation frame [10] structured as a primary activity (i.e., multi-entity involved relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates.
+Each role explicitly defines the entity’s participation in the activity, e.g., <agent>person</agent><box>[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]</box>.
 3.2 Supervised Fine-Tuning
-As mentioned above, directly adopting GRPO faces challenges in generating format-inconsistent scene graph captions and grounded situation frame outputs.
-To this end, we adopt supervised fine-tuning that serves as a foundational stage for task-specific format alignment (c.f., Figure 2(a)).
-However, utilizing merely the answer format as the SFT target risks compromising the model’s reasoning capability by encouraging rigid adherence to syntactic structures [3].
-To mitigate this, we devise the cognitive Chain-of-Thoughts (CoT), which incorporates stepwise cognitive processes such as object classification, object grounding, visual relationship inference, etc.
-This ensures the model retains its ability to perform multi-step reasoning while adhering to output formats.
-To avoid the MLLM overfitting to specific CoT patterns during SFT, we devise a cognitive CoT generation prompt that explicitly guides an extra strong MLLM (e.g., Qwen 2.5-VL [31]) to generate step-by-step reasoning processes.
-To be specific, this cognitive CoT generation prompt consists of task definition, scene graph caption/grounded situation frame examples, and GT scene graph caption/grounded situation frame of the current image.
-The former two make the MLLM generate the correct task format.
-Meanwhile, the latter one anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors.
-The detailed Cognitive CoT generation prompt is left in the Appendix.
-These cognitive CoT are explicitly encapsulated within <think> tags, during SFT.
-By leveraging this structured training signal, the MLLM learns to generate format-compliant outputs and grounded reasoning processes as an initial capability, which is critical for subsequent CoT-guided generalization under the GRPO.
+As discussed in §1, directly applying GRPO encounters challenges in generating consistently formatted scene graph captions and grounded situation frame outputs.
+To address this, we employ SFT as a foundational stage to establish task-specific output format adherence (c.f., Figure 2(a)).
+However, relying solely on the target answer format during SFT risks compromising the model’s underlying reasoning capabilities by encouraging a superficial alignment with formulated structures [13].
+To mitigate this potential issue, we introduce the cognitive chain-of-thought, a methodology that integrates stepwise cognitive processes such as object classification, object grounding, and visual relationship inference.
+These cognitive CoTs are explicitly encapsulated within <think> tags, during SFT.
+This ensures that the model retains its capacity for multi-step reasoning while learning to produce structured outputs conforming to the desired format.
+To streamline the think process and prevent the MLLM from overfitting to specific CoT patterns during SFT, we adopt a specific-to-general progressive training strategy:
+Template-based CoT.
+For the specific learning phase, we devise a fixed, template-based CoT.
+As illustrated in Figure 2(a), this template comprises a predefined sequence of steps for both binary and N-ary relation detection.
+To elaborate, for binary relation detection, the template includes: 1) Object Existence (e.g., the objects present ...); 2) Object Localization (e.g., the objects are located ...); and 3) Relation Existence (e.g., the relations present ...).
+For N-ary relation detection, it involves: 1) Activity Recognition (e.g., the primary activity is ...); 2) Entities and Roles Recognition (e.g., the entities engaged in the activity are...); and 3) Entity Localization (e.g., the entities are located in ...).
+MLLM-generated CoT.
+For the general reasoning process, we design a cognitive CoT generation prompt that explicitly makes an extra strong MLLM (e.g., Qwen 2.5-VL [39]) to generate step-by-step reasoning processes.
+To be specific, this cognitive CoT generation prompt consists of a task definition, a ground-truth scene graph of the current image, and a CoT generation instruction.
+The former one makes the MLLM generate the correct task format.
+Meanwhile, the latter two anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors.
+The detailed cognitive CoT generation prompt is left in the appendix §B.
+Progressive CoT Guidance.
+With a progressive paradigm that involves initial training on easy-to-learn fully template-based CoTs and subsequent fine-tuning on a small number of MLLM-generated CoTs, the MLLM gains the foundational capability to generate format-compliant outputs and grounded reasoning processes, which is critical for GRPO training.
 3.3 Group Relative Policy Optimization
-Following the Deepseek-R1 [7], the Group Relative Policy Optimization (GRPO) is adopted as the reinforcement learning (RL) algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b).
+Following the Deepseek-R1 [12], the group relative policy optimization is adopted as the RL algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b).
 Distinct from critic-dependent methods [40], GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity [12].
-Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{i}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge.
-The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging following objectives:
-$J_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}$ (1)
-$\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$
+Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{1},o_{2},\ldots,o_{G}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge.
+The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging the following objectives:
+$J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$ (1)
 where $\rho_{i}$ = $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ quantifies policy change, $\epsilon$ controls clipping thresholds, and $\beta$ penalizes deviations via the KL divergence term.
 The advantage score $A_{i}$ is standardized as:
-$A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
-here, $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
-$D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1,$ (3)
-which ensures controlled exploration without excessive divergence from $\pi_{\theta_{\text{ref}}}$.
-The reward function $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding.
+$\small A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
+here $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
+$\small D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1.$ (3)
+This term ensures controlled exploration without excessive divergence from $\pi_{\theta_{\text{ref}}}$.
+The reward $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding.
 It incorporates rule-based rewards, which consist of three crucial components: Format compliance rewards, binary relation rewards, and N-nary relation rewards.
@@
 Format Rewards. This reward enforces strict adherence to the reasoning template format: <think> thinking process </think><answer> answer </answer>. The reward function is defined as:
-$r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
+$\small r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
 This binary mechanism ensures syntactic compliance while allowing content flexibility.
 Binary Relation Rewards.
-Inspired by established evaluation practices [4, 1], we devise the sample-level triplet Recall and sample-level mean Recall (mRecall) for each predicate category in the response of each sample.
+Inspired by established evaluation practices [17, 8], we devise the sample-level triplet recall $R$ and sample-level mean recall $mR$ in the response of each sample.
+The former is the ratio of correctly predicted triplets to the total ground-truth triplets in a sample; the latter is the average recall across all predicate categories.
 Specifically, a triplet is deemed correct if:
-  • Triplet Accuracy: All entity categories (subject, predicate, object) in this sample are correctly identified.
+  • Triplet Accuracy: All entity and relation categories (subject, predicate, object) in this sample are correctly identified.
   • Spatial Localization: The predicted bounding boxes for the subject and object achieve an Intersection over Union (IoU) >= 0.5 with their respective ground-truth boxes.
 The final binary relation reward is computed as a weighted combination of these two metrics:
-$r_{\text{binary}}(o)=\alpha\cdot Recall+(1-\alpha)\cdot mRecall,$ (5)
+$\small r_{\text{binary}}(o)=\alpha\cdot R+(1-\alpha)\cdot mR,$ (5)
 where $\alpha$ balances the trade-off between two metrics.
 N-ary Relation Rewards.
-Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary Relation Rewards, which assess the model’s ability to detect primary activity and multi-entity interactions.
-Specifically, we calculate the verb accuracy $Acc_{v}$ and noun accuracy $Acc_{n}$:
-  • Verb Accuracy: The predicted primary activity (verb) in the grounded situation frame is the same as the ground-truth activity.
-  • Noun Accuracy: All entities involved in the event must satisfy IoU >=0.5 with their ground-truth bounding boxes, and their roles must align with the event’s semantic structure.
+Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary relation rewards, which assess the model’s ability to detect primary activity (verb) and multi-entity interactions.
+Specifically, we calculate the entity value $V_{e}$ and grounded value $V_{grnd}$ for each sample [10].
+These respectively represent the correct proportion in predicted entity categories and roles, and the spatial localization of those entities:
+  • Entity Accuracy: An entity within an event is considered correct if its predicted category and semantic role align with the ground truth, and the predicted activity matches the ground-truth.
+  • Spatial Localization: The predicted bounding box of each role obtains IoU >=0.5 with the ground-truth’s bounding box.
 The eventual N-ary relation reward is computed as a weighted combination of these two accuracies:
-$r_{\text{n-ary}}(o)=\beta\cdot Acc_{v}+(1-\beta)\cdot Acc_{n},$ (6)
-where hyperparameter $\beta$ balances the trade-off between event type precision and multi-entity structural accuracy.
+$\small r_{\text{n-ary}}(o)=\beta\cdot V+(1-\beta)\cdot V_{grnd},$ (6)
+where the hyperparameter $\beta$ balances the trade-off between the two values.
+As the reward’s calculation depends on accurate verb prediction, they jointly evaluate primary activity and multi-entity interactions.
 During training, we design a multi-task gate that dynamically selects task-specific rewards based on the presence of <ref> tags in the solutions.
@@
 4.1 Experimental Setup
-4.1.1 Datasets
-PSG.
-The Panoptic Scene Graph (PSG) dataset [4] is built upon the COCO dataset [33], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing.
+Datasets.
+1) Panoptic Scene Graph (PSG): The official PSG dataset [17], built upon the COCO dataset [41], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing.
 It integrates 80 “thing” object categories and 53 “stuff” object categories, alongside 56 relation categories.
-To enable open-ended scene graph generation, we followed ASMv2 [1], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training.
-SWiG.
-The SWiG dataset [5] extends the imSitu dataset [34] by adding extra bounding box annotations for visible semantic roles, with 69.3% of semantic roles localized.
+Notably, ASMv2 [8] reconstructs the dataset into a scene graph caption format, utilizing a split of 42,250 training images and 1,000 evaluation images.
+To mark fair comparison, we followed ASMv2 [8], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training.
+Besides, we also present results using the conventional scene graph format, i.e., a list of visual triplets structured as [subject, subject bounding box, object, object bounding box, relation] on the official dataset [^1] More details of the scene graph format are in the appendix §C..
+2) SWiG: The SWiG dataset [10] extends the imSitu dataset [42] by adding extra bounding box annotations for semantic roles, with 69.3% of semantic roles localized.
 Each image is annotated with a verb and a variable number of semantic roles (1$\sim$6 roles per image), where each verb is associated with three verb frames annotated by independent human annotators.
@@
 Unlike traditional output structures that merely predict categories and bounding boxes, we constructed grounded situation frames by filling verb-centric templates (e.g., “The AGENT drinks a LIQUID from a CONTAINER”) with three critical components: role tags (e.g., <agent>, entity categories (e.g., child), and coordinates of corresponding bounding boxes in <box> tags.
-4.1.2 Evaluation Metrics
-Binary Relation Detection.
-Following the standard SGG evaluation protocol [15, 1], we reported 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object.
+Evaluation Metrics.
+For binary relation detection, we followed the standard SGG evaluation protocol [24, 8] 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object.
 2) mean Recall (mRecall): The average recall across all predicate categories.
-N-ary Relation Detection.
-We adopted the same five evaluation metrics as [5]: 1) verb: the accuracy of verb prediction.
+As for N-ary relation detection, we adopted the same five evaluation metrics as [10]: 1) verb: the accuracy of verb prediction.
 2) value: the noun accuracy for individual semantic role.
@@
 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
-4.1.3 Implementation Details
-We utilized Qwen2.5-VL-3B [31] as our base model.
-Relation-R1 was trained on an 8xA100 GPU server.
-During the SFT phase, we employed a total batch size of 64, with 8 samples processed per training step.
-For the GRPO component, we used a total batch size of 32.
-The learning rates were set to 1e-5 for SFT and 1e-7 for GRPO, respectively.
-Table 1: Performance(%) comparison with various SGG models on the PSG dataset [4].
-            Method Model Size Answer Recall mRecall Mean
-IMP [35] ${}_{\text{CVPR'17}}$ Close-ended 16.50 6.50 11.50
+Implementation Details. Due to space constraints, details are provided in the appendix §A.
+[…]
+         Method Model Size Format Recall mRecall Mean
+Close-ended
+IMP [16] ${}_{\text{CVPR'17}}$ 16.50 6.50 11.50
 MOTIFS [25] ${}_{\text{CVPR'18}}$ 20.00 9.10 14.55
@@
 PSGFormer [17] ${}_{\text{ECCV'22}}$ 18.60 16.70 17.65
-TextPSG [38] ${}_{\text{ICCV'23}}$ Open-ended 4.80 - -
-ASMv2 [1] ${}_{\text{ECCV'24}}$ 13B 14.20 10.30 12.23
+Open-ended
+TextPSG [45] ${}_{\text{ICCV'23}}$ Standard Scene Graph 4.80 - -
+R1-SGG [46] ${}_{\text{ArXiv'25}}$ 2B 27.83 17.03 22.43
+R1-SGG [46] ${}_{\text{ArXiv'25}}$ 7B 28.77 17.55 23.16
+Relation-R1 (Ours) 3B 25.87 21.32 23.60
+ASMv2 [8] ${}_{\text{ECCV'24}}$ 13B Scene Graph Caption 14.20 10.30 12.23
 LLaVA-SpaceSGG [9] ${}_{\text{WACV'25}}$ 13B 15.43 13.23 14.33
-Relation-R1 (Ours) 3B 19.59 18.30 18.95
+Relation-R1 (Ours) 3B 22.33 20.07 21.20
 Table 2: Performance(%) comparison with various GSR models on the SWiG [10] dataset.
-         Method Answer Verb Value Value-all Gnd Value Gnd Value-all
+         Method Answer Verb Value Value-all Grnd Grnd-all
 ISL [10] ${}_{\text{ECCV'20}}$ Close-ended 39.36 30.09 18.62 22.73 7.72
@@
 OpenSU [47] ${}_{\text{ICCVW'23}}$ Open-ended 50.10 41.20 26.56 34.27 15.70
-Relation-R1 (Ours) 53.76 43.55 28.23 36.79 26.84
-Table 3: Few-Shot GSR performance comparison. $\dagger$ denotes employing the SFT phrase without CoT.
-         Method Setting Verb Value Value-all Gnd Value Gnd Value-all
+Relation-R1 (Ours) 57.26 46.66 30.92 40.21 30.18
+4.2 Comparison with State-of-the-Art Methods
+We compared our Relation-R1 for unified relation comprehension with the existing state-of-the-art binary relation detection and N-ary relation detection approaches in Table 1 and Table 2, respectively.
+[…]
+[…]
+4.3 Diagnostic Experiment
+To gain more insights and reduce the mutual influence between the two tasks during training, we conducted a set of comprehensive studies on the Relation-R1 trained with binary and N-ary relation detection, respectively.
+Table 3: Performance comparison (%) across various cognitive CoT strategies trained on single binary relation detection and $N$-ary relation detection, separately.
+Blue indicates metrics without correct verb constraints.
+  Binary Relation $N$-ary Relation
+Method CoT Recall mRecall Verb Value Value-all Grnd Grnd-all
+SFT - 14.83 13.86 56.64 42.65 (60.45) 22.11 (25.21) 37.70 (54.05) 16.35 (19.02)
+SFT + RL Template-based 20.24 17.31 58.38 47.75 (66.74) 32.00 (35.67) 41.27 (55.88) 31.16 (38.29)
+SFT + RL MLLM-generated 20.66 20.30 53.00 42.27 (63.86) 26.67 (30.79) 35.69 (52.43) 25.89 (33.95)
+SFT + RL Progressive 22.57 20.57 71.04 61.26 (78.19) 44.98 (49.73) 49.43 (61.25) 36.09 (42.35)
+Cognitive Chain-of-Thoughts.
+Table 3 compares different cognitive CoT strategies for binary and N-ary relation detection.
+As seen, introducing RL with various CoT strategies (SFT + RL) demonstrates clear benefits.
+The template-based CoT improves binary relation Recall to 20.24% and mRecall to 17.31%.
+For N-ary relations, it enhances performance across the board compared to the constrained SFT baseline, for instance, achieving 47.75% on Value and 31.16% on Grnd-all.
+The MLLM-generated CoT slightly outperforms the template-based approach in binary relation detection, but underperforms in N-ary relation detection.
+This is because N-ary relation detection metrics are often limited by verb accuracy, the model may generate a verb different from the GT yet still semantically plausible.
+Notably, when the verb constraint is removed, MLLM-generated CoT yields significant gains, particularly in Grnd-all (+14.93%).
+While template-based and MLLM-generated CoTs offer improvements over SFT, the progressive strategy stands out.
+This approach significantly outperforms all others across all metrics, achieving 22.57% Recall in binary relation detection and top scores (e.g., 71.04% Verb accuracy) in N-ary relation detection.
+These results highlight the superior ability of the progressive CoT to guide complex relation detection.
+Figure 3: The statistics of binary relation rewards.
+Figure 4: The statistics of $N$-ary relation rewards.
+Reward Analysis.
+We visualized the reward variation curves for binary and N-ary relation detection in Figure 3 and Figure 4, respectively.
+It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT.
+With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend.
+In binary relation detection, guided with progressive CoT yields a slightly enhanced overall reward trajectory, although the improvement is not significant.
+This might be since the gap between the CoT generated by MLLM (e.g., Qwen [39]) and that produced by the template is not obvious.
+Conversely, N-ary relation detection exhibits a significant surge in reward, primarily attributed to the model’s emergent capability to generate synonymous grounded situation frames (c.f.
+§5).
+Completion Length Analysis.
+Figure 5: The statistics of binary relation completion lengths.
+Figure 6: The statistics of $N$-ary relation completion lengths.
+The evolution of completion length during GRPO training is depicted in Figure 5 and Figure 6.
+Consistent with the reward trends, the completion lengths for binary relation detection exhibit a gradual increase as GRPO progresses, suggesting the generation of more detailed and descriptive outputs over time.
+Notably, the template-based CoT strategy yields significantly shorter completions compared to the more flexible MLLM-generated and progressive CoT strategies, since only the cognitive process is involved.
+The progressive CoT guidance, by enabling the model to explore synonymous relational expressions and incorporate more nuanced details, leads to an increase in the average completion length.
+Table 4: Few-Shot GSR performance comparison.
+         Method Setting Verb Value Value-all Grnd Grnd-all
 GSRTR [11] ${}_{\text{BMVC'21}}$ One-Shot 0.35 0.10 0.01 0.08 0.00
 CoFormer [36] ${}_{\text{CVPR'22}}$ 0.63 0.23 0.02 0.14 0.01
-Relation-R1 (Ours) 27.31 10.21 1.43 9.35 1.22
 Relation-R1 (Ours) 31.67 15.37 4.15 14.01 3.41
+$w$/$o$ RL 27.31 10.21 1.43 9.35 1.22
 GSRTR [11] ${}_{\text{BMVC'21}}$ 4-Shot 6.95 4.32 1.83 2.79 0.65
 CoFormer [36] ${}_{\text{CVPR'22}}$ 12.39 8.27 3.93 5.28 1.25
-Relation-R1 (Ours) 34.88 17.14 3.94 15.63 3.38
 Relation-R1 (Ours) 38.02 21.87 7.39 19.64 5.92
-(a) SGG.
-(b) GSR.
-Figure 3: The statistics of rewards during training.
-Figure 4: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG. The orange area denotes the predicate distribution of the training set.
-Figure 5: Qualitative results of binary relation detection (scene graph generation).
-Figure 6: Qualitative results of N-ary relation detection (grounded situation recognition).
-4.2 Quantitative Comparison Result
-4.2.1 Binary Relation Detection
-Open-ended SGG.
-We evaluated Relation-R1 on PSG, comparing it against various established and recent SGG models, as detailed in Table 1.
-Positioned within the open-ended generation category, Relation-R1 achieved SOTA performance with the highest mRecall of 18.30% and overall Mean of 18.95%, substantially surpassing prior open-ended methods like LLaVA-SpaceSGG (13.23% mRecall, 14.33% Mean).
-It also achieved a competitive Recall of 19.59%.
-Notably, Relation-R1 attained these leading results using a significantly smaller 3B parameter model compared to the 13B models used by recent competitors like ASMv2 and LLaVA-SpaceSGG.
-In addition, we visualized predicate-level comparison against ASMv2 and the predicate distribution in Figure 4.
+$w$/$o$ RL 34.88 17.14 3.94 15.63 3.38
+Few-Shot N-ary Relation Detection.
+We also evaluated Relation-R1’s few-shot generalization capability (Table 4).
+In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Grnd-all) compared to CoFormer’s [36] 0.63% and 0.01%, respectively.
+This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Grnd-all), significantly surpassing previous methods.
+These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
+Joint Binary and N-ary Relation Detection.
+Table 5: Performance comparison (%) across various cognitive CoT strategies trained on joint binary relation detection and $N$-ary relation detection, separately.
+Blue indicates metrics without correct verb constraints.
+  Binary Relation $N$-ary Relation
+Method CoT Recall mRecall Verb Value Value-all Grnd Grnd-all
+SFT - 14.89 12.66 56.23 45.81 (66.00) 30.31 (34.06) 38.59 (54.11) 28.77 (36.18)
+SFT + RL Template-based 22.33 20.07 57.26 46.66 (66.14) 30.92 (34.75) 40.21 (55.50) 30.18 (37.64)
+SFT + RL MLLM-generated 19.59 18.30 53.00 42.27 (65.11) 26.67 (34.65) 35.69 (55.36) 25.89 (37.61)
+SFT + RL Progressive 27.19 26.12 73.25 66.18 (82.54) 52.90 (57.83) 55.20 (67.13) 42.36 (48.81)
+Beyond training Relation-R1 solely on binary or N-ary relation detection tasks (as discussed in §3), we also investigated the joint training of both tasks within the same Relation-R1 model.
+The results of this joint training are presented in Table 5333The Grnd-all metric excludes ground-truth bounding boxes represented as [-1, -1, -1, -1] for calculation..
+From the analysis of these results, three primary conclusions can be drawn: 1) When utilizing both the SFT and RL strategies, our Relation-R1 consistently achieves higher performance across nearly all metrics for both binary and N-ary relation detection tasks.
+Similar to single-task training, the MLLM-generated CoT strategy underperforms in the N-ary task.
+This is likely due to the inherent generality of the CoT generated by Qwen [39], which may not precisely align with the ground-truth annotations, consequently leading to lower Verb accuracy.
+2) Our proposed progressive CoT-guided paradigm consistently achieves the highest performance across all evaluated metrics for both tasks in the joint learning setting.
+3) The joint progressive CoT-guided training across both binary and N-ary tasks fosters a synergistic mutual promotion effect.
+This is evidenced by the notable gains observed: for instance, a Recall of 27.19% in Table 5 vs.
+22.57% in Table 3 (for binary relation detection), and a Verb accuracy of 73.25% in Table 5 vs.
+71.04% in Table 3 (for N-ary relation detection).
+5 Qualitative Results
+Figure 7: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG [17] dataset. The orange area denotes the predicate distribution of the training set.
+Figure 8: Qualitative results of binary relation detection.
+Figure 9: Qualitative results of N-ary relation detection.
+Figure 10: Qualitative results of synonymous in N-ary relation detection.
+Comparison over All Predicates.
+We visualized predicate-level comparison against ASMv2 [8] and the predicate distribution in Figure 7.
 As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates.
 These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
-4.2.2 N-ary Relation Detection
-Open-ended GSR.
-We evaluated Relation-R1 for N-ary relation detection on SWiG, comparing it against prior work in Table 2.
-Within the open-ended setting, Relation-R1 established a new SOTA, achieving top performance across all reported metrics: Verb (53.76%), Value (43.55%), Value-all (28.23%), Gnd Value (36.79%), and Gnd Value-all (26.84%).
-Notably, Relation-R1 significantly surpassed the previous best open-ended model, OpenSU [39], exhibiting substantial gains, particularly with a remarkable +11.14% absolute improvement on the challenging Gnd Value-all metric.
-This clearly demonstrated Relation-R1’s superior capacity for comprehending and grounding complex N-ary relationships in visual scenes.
-Few-Shot GSR.
-We also evaluated Relation-R1’s few-shot generalization capability (Table 3).
-In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Gnd Value-all) compared to CoFormer’s [28] 0.63% and 0.01%, respectively.
-This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Gnd Value-all), significantly surpassing previous methods.
-These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
-4.3 Qualitative Result
-Figure 5 and Figure 6 provided qualitative illustrations of Relation-R1.
-For Binary Relation Detection (Figure 5), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference.
-For instance, after identifying the central bed, Relation-R1 pinpoints related objects like the rug and chair and explicitly grounds the inferred spatial relationships (on, beside) between them.
-This systematic approach demonstrates the successful decomposition of the scene into accurate, grounded object-relation pairs.
-Similarly, for N-ary Relation Detection (Figure 6), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box.
+Binary and N-ary Relation Detection Results.
+We visualized some examples of binary and N-ary relation detection in Figure 8 and Figure 9, respectively.
+As seen, for binary relation detection (c.f.
+Figure 8), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference.
+For instance, Relation-R1 can first focus on the main objects (e.g., “cat” and “train”) and then infer their positions and relationships.
+Similarly, for N-ary relation detection (c.f.
+Figure 9), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box.
 This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
-4.4 Diagnostic Experiment
-To gain more insights into Relation-R1, we conducted a set of comprehensive studies on SGG and GSR.
-Reward Statistics.
-We visualized the variation curves of rewards in the SGG task and the GSR task, respectively, as displayed in Figure 3(a) and Figure 3(b).
-It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT.
-With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend.
+Synonymous Visualization.
+Figure 10 displays synonymous results observed in N-ary relation detection.
+As evinced by these results, our progressive CoT-guided training paradigm empowers the model to generate diverse synonymous expressions, encompassing entities, semantic roles, and activities.
+For instance, in Figure 10(top), Relation-R1 demonstrates the ability to output synonymous agents such as “person” and “man”.
+Furthermore, the model can generate semantically similar and plausible activities, exemplified by “barbecues” and “grills” (c.f.
+Figure 10(bottom)).
+Even within the same activity “rests”, it can differentiate and utilize varying semantic role names like “item” and “bodypart” (c.f.
+Figure 10(row 2)).
+These synonymous outputs indicate a more flexible and robust semantic understanding beyond rigid lexical matching, thereby significantly enhancing the model’s generalization capabilities to diverse linguistic formulations and real-world variations.
 6 Conclusion
-This paper reveals critical limitations of existing MLLMs in visual relation understanding, particularly concerning N-ary relationships and the detrimental effects of language-induced bias.
-We propose Relation-R1, an advanced, unified framework designed to jointly tackle both binary and N-ary relational comprehension.
-Our core innovation lies in a two-stage approach: first establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, and subsequently refining the model’s outputs using RL with GRPO.
-This RL stage, guided by carefully designed multi-rewards, explicitly promotes visual-semantic grounding and robust generalization.
-Extensive experiments conducted on the challenging PSG and SWiG datasets validate the effectiveness of Relation-R1, demonstrating SOTA performance across both binary and N-ary relation detection tasks.
-By successfully integrating CoT-guided SFT with R1-inspired RL refinement, Relation-R1 represents a solid step towards equipping MLLMs with more human-like relational reasoning capabilities.
+This paper reveals critical limitations in MLLM visual relation understanding, particularly for N-ary relationships and language-induced bias.
+We propose Relation-R1, a unified framework with a two-stage approach: establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, followed by RL with GRPO for refinement, promoting visual-semantic grounding and robust generalization.
+Furthermore, progressively applied cognitive CoT guidance, transitioning from template-based to MLLM-generated reasoning, further enhances generalization, especially for synonymous N-ary relations.
+Extensive experiments on PSG and SWiG validate Relation-R1’s effectiveness, demonstrating SOTA performance in both binary and N-ary relation detection.
+Integrating progressive CoT-guided SFT with R1-inspired RL refinement, Relation-R1 advances MLLMs towards more human-like relational reasoning.
+Appendix A Implementation Details
+We utilized Qwen2.5-VL-3B [39] as our base model.
+Relation-R1 was trained on an 8xA100 80GB GPU server.
+During the SFT phase, we employed a total batch size of 32, with 4 samples processed per training step.
+For the GRPO component, we used a total batch size of 48.
+The learning rates were set to 2e-5 for SFT and 1e-6 for GRPO, respectively.
+Qwen2.5-VL-72B [39] served as the CoT generator for producing MLLM-generated CoTs.
+Progressive training began with supervised fine-tuning on binary and N-ary relation detection tasks for 2 epochs using template-based CoT, followed by fine-tuning on 4k MLLM-generated CoT samples.
+After that, GRPO RL training was conducted: 2.4k steps on the N-ary relation detection task, 3.6k steps on the binary relation detection task, and finally 2k steps on both tasks.
+Appendix B Cognitive CoT Generation Prompt
+Question: Scene graph caption is a textual description of an image that incorporates references for objects and their relationships: 1.
+Use <ref>object class</ref><box>[[x1,y1,x2,y2],…,[x1,y1,x2,y2]]</box> to denote all subject and object regions of a relation triplet in the caption.
+2.
+Use <pred>predicate class</pred><box>subject coordinates list</box><box>object coordinates list</box> to denote all relation triplet regions.
+3.
+Each sentence must have at least one relation triplet region.
+Given the image, a reasonable scene graph caption is: {ground-truth caption} To make you generate a reasonable scene graph caption, please give the step-by-step reasoning process in <think> </think> tags.
+Figure S1: Scene graph caption CoT generation prompt.
+Question: Grounded situation frame is a multimodal description combining textual semantics with visual localization information, where key situational elements (agent, action, object, place) are associated with both semantic role labels and corresponding bounding box coordinates in the image.
+It follows the pattern: <role>entity class</role><box>[x1,y1,x2,y2]</box>.
+Given the image, a reasonable grounded situation frame is: {ground-truth frame} To make you generate a reasonable grounded situation frame, please give the step-by-step reasoning process in <think> </think> tags.
+Figure S2: Grounded situation frame CoT generation prompt.
+As mentioned in §3.2, we employ two distinct prompts (Figure S1 and Figure S2) to guide cognitive CoT generation for binary and N-ary relation detection, respectively.
+Each prompt strategically comprises: 1) a concise task definition outlining the detection objective; 2) an illustrative in-context example demonstrating the desired reasoning steps and output format; and 3) a specific instruction prompting the model to generate CoT of the current image.
+Appendix C Scene Graph Format
+Question: Generate a scene graph for the provided image as a structured list containing [subject entity, [x1, y1, x2, y2] bounding box coordinates of subject, object entity, [x1, y1, x2, y2] bounding box coordinates of object, relationships between subject and object].
+First output the thinking process in <think> </think> tags and then output the final answer in <answer> </answer> tags.
+Figure S3: Scene graph format prompt used for training.
+The scene graph format prompt employed for GRPO training is displayed in Figure S3.
+Furthermore, to enable the model’s generation of think process, we utilize templated-CoT (c.f.
+§3.2) for supervised fine-tuning.
+Appendix D Discussions
+Limitations.
+While Relation-R1 demonstrates state-of-the-art performance in visual relation understanding, several limitations warrant discussion and offer avenues for future research.
+The first one is computational cost.
+Like other GRPO-based frameworks, the proposed framework introduces a multi-stage training process that requires substantial time to collect a training batch.
+The second one is the interpretability of GRPO refinements.
+While CoT aims to make the initial reasoning process more transparent, the subsequent policy optimization by GRPO, while guided by multi-rewards, might still operate as a “black box” to some extent.
+Broader Impacts.
+The development of Relation-R1 promises significant positive impacts, such as more reliable AI systems with reduced hallucinations for applications in robotics and content analysis, enhanced accessibility tools for the visually impaired through richer scene descriptions, and advanced information retrieval based on complex multi-entity interactions.
+However, ethical considerations include the potential misuse for generating sophisticated misinformation, the risk of amplifying societal biases present in training data, and heightened surveillance capabilities if deployed without oversight.

### v3

@@
-Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Kwang-Ting Cheng1,2  Long Chen1†
+Relation-R1: Progressively Cognitive Chain-of-Thought Guided Reinforcement Learning for Unified Relation Comprehension
 Abstract
@@
 However, they remain limited in visual relation understanding, struggling even with binary relation detection, let alone N-ary relations involving multiple semantic roles.
-The core reason is the lack of modeling for structural semantic dependencies among multi-entities, leading to unreliable outputs, hallucinations, and over-reliance on language priors (e.g., defaulting to “person drinks a milk” if a person is merely holding it).
+The core reason is the lack of modeling for structural semantic dependencies among multi-entities, leading to over-reliance on language priors (e.g., defaulting to “person drinks a milk” if a person is merely holding it).
 To this end, we propose Relation-R1, the first unified relation comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided supervised fine-tuning (SFT) and group relative policy optimization (GRPO) within a reinforcement learning (RL) paradigm.
@@
 Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
-1 Introduction
-Recent advancements in MLLMs have significantly enhanced holistic image understanding [1, 2] and object-level grounding [3, 4, 5, 6, 7] capabilities, enabling tasks like region captioning and referring question answering.
-However, current MLLMs exhibit critical limitations in relational scene understanding — a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
+Codegithub.com/HKUST-LongGroup/Relation-R1
+Introduction
+Recent advancements in MLLMs have significantly enhanced holistic image understanding (liu2023visual; zhu2024minigpt) and object-level grounding (youferret; zhang2024llava; rasheed2024glamm; yuan2024sr; yuan2025voxel; lai2024lisa; peng2024kosmos) capabilities, enabling tasks like region captioning and referring question answering.
+However, current MLLMs exhibit critical limitations in relational scene understanding – a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
 Figure 1: Illustrations of our Relation-R1 and other MLLMs.
@@
 (b) Relation Comprehension MLLMs: Model pairwise object relations but lack generalization to N-ary interactions.
-(c) Unified Relation Comprehension MLLMs: Jointly handle binary and $N$-ary relation detection.
-Despite notable progress in binary relation detection, i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), current models [8, 9] struggle to capture N-nary relationships.
-Accurately modeling such relationships requires models to identify multiple N-th entities engaged in the relation with their distinct semantic roles (what, who, and where concepts) [10, 11].
+(c) Unified Relation Comprehension MLLMs: Jointly handle binary and NN-ary relation detection.
+Despite notable progress in binary relation detection, i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), current models (wang2024all; llava_spacesgg2025) struggle to capture N-ary relationships.
+Accurately modeling such relationships requires models to identify multiple N-th entities engaged in the relation with their distinct semantic roles (what, who, and where concepts) (pratt2020grounded; cho2021grounded).
 The neglect of these crucial structural semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, may result in suboptimal relation triplets, e.g., “child-drinking-glass”.
-Furthermore, these unreliable and superficial pair-wise relations may make the MLLMs prone to both hallucinations and over-reliance on language priors rather than visual-semantic cues (e.g., evidence from semantically grounded visual regions).
+Furthermore, these unreliable and superficial pair-wise relations may make the MLLMs prone to over-reliance on language priors rather than visual-semantic cues (e.g., evidence from semantically grounded visual regions).
 Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings.
@@
 For example, a model might struggle to recognize “drinking” from “milk” if its language training strongly associates this action with “glass”.
-Fortunately, DeepSeek-R1-Zero [12] has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL.
-However, due to the inherent difficulty of reference models in understanding structural relationships, the direct application of RL often struggles to produce consistently formatted outputs that are claimed in the input prompt [13].
+Fortunately, DeepSeek-R1-Zero (DeepseekR1) has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL.
+However, due to the inherent difficulty of reference models in understanding structural relationships, directly applying RL often struggles to produce consistently formatted outputs that are claimed in the input prompt (chu2025sft).
 Conversely, when solely adopting supervised fine-tuning, the model suffers from poor generalization due to its overfitting to fixed training patterns and the limited diversity of annotated relational data.
-To this end, inspired by the breakthrough of R1-series [12, 14, 15] in enabling reasoning and the ability of SFT to provide structured guidance, we propose a unified two-stage relation comprehension framework, Relation-R1, which combines the advantages of both supervised fine-tuning and reinforcement learning to empower MLLMs with relational reasoning and generalization capability:
+To this end, inspired by the breakthrough of R1-series (DeepseekR1; yang2025r1onevision; zhao2025r1) in enabling reasoning and the ability of SFT to provide structured guidance, we propose a unified two-stage relation comprehension framework, Relation-R1, which combines the advantages of both supervised fine-tuning and reinforcement learning to empower MLLMs with relational reasoning and generalization capability:
 Stage 1.
@@
 RL: Refines the structured outputs generated by the SFT model through group relative policy optimization.
-The learning process is guided by multi-rewards: 1) format rewards, which incentivize the generation of thinking process [12]; 2) binary relation rewards, which encourage accurate identification of pairwise relationships between entities; and 3) N-ary relation rewards, which promote the comprehensive understanding of more complex, multi-entity activity.
+The learning process is guided by multi-rewards: 1) format rewards, which incentivize the generation of thinking process (DeepseekR1); 2) binary relation rewards, which encourage accurate identification of pairwise relationships between entities; and 3) N-ary relation rewards, which promote the comprehensive understanding of more complex, multi-entity activity.
 This stage employs policy gradient updates to cultivate the model to explore various potential solutions and optimizes its output based on defined multi-rewards, fostering robust relational reasoning and generalization.
@@
 Especially in the N-ary relations, the capacity to explore synonymous relational expressions has emerged.
-We validate the effectiveness of Relation-R1 on both the binary relation comprehension task (e.g., scene graph generation [16, 8]) and N-ary relation comprehension task (e.g., grounded situation recognition [10]) across diverse experimental settings.
-Experiments on the PSG [17] dataset and SWiG [10] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts.
+We validate the effectiveness of Relation-R1 on both binary relation comprehension (e.g., scene graph generation (xu2017scene; wang2024all)) and N-ary relation comprehension task (e.g., grounded situation recognition (pratt2020grounded)) across diverse experimental settings.
+Experiments on the PSG (yang2022panoptic) and SWiG (pratt2020grounded) datasets demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts.
 In summary, our contributions are fourfold:
@@
   • We comparatively analyze template-based and MLLM-generated CoTs, demonstrating that progressively guiding the learning process with CoTs in a specific-to-general manner enhances generalization capabilities, particularly in capturing synonymous N-ary relations.
-  • Experimental results on the prevalent relation understanding datasets (e.g., +6.84$\sim$6.90gains on PSG dataset) under different settings demonstrate the effectiveness of Relation-R1.
-2 Related Work
+[…]
+Related Work
 Reasoning in MLLMs.
-Recent LLMs have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [18, 12].
-A pivotal breakthrough, DeepSeek-R1 [12], leverages large-scale RL with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains.
-MLLMs extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [14, 19, 20] For multimodal mathematical reasoning, frameworks such as R1-Onevision [14], Virgo [19], and MM-Eureka [21] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [20] combines cold-start initialization with GRPO to refine complex CoT reasoning.
-For fine-grained classification and grounding tasks, Visual-RFT [22] and R1-Omni [15] utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning.
-As for the pixel-level understanding task, SegZero [23] employs reinforcement learning to achieve high-resolution semantic segmentation.
+Recent LLMs have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks (jaech2024openai; DeepseekR1).
+A pivotal breakthrough, DeepSeek-R1 (DeepseekR1), leverages large-scale RL with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains.
+MLLMs extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards (yang2025r1onevision; du2025virgo; huang2025vision; shen2025tarpro) For multimodal mathematical reasoning, frameworks such as R1-Onevision (yang2025r1onevision), Virgo (du2025virgo), and MM-Eureka (meng2025mm) jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 (huang2025vision) combines cold-start initialization with GRPO to refine complex CoT reasoning.
+For fine-grained classification and grounding tasks, Visual-RFT (liu2025visual) and R1-Omni (zhao2025r1) utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning.
+As for the pixel-level understanding task, SegZero (liu2025seg) employs reinforcement learning to achieve high-resolution semantic segmentation.
 Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts.
@@
 Scene Graph Generation (SGG).
-SGG is a critical task in scene understanding, with prior approaches falling into two categories: 1) Two-stage SGG, which sequentially detects objects and infers pairwise relations but suffers from error propagation [24, 25, 26, 27]; 2) One-stage SGG, which unifies detection and relation prediction in end-to-end frameworks (e.g., DETR-based methods [28, 29, 30]) but often lacks multi-role involved relation modeling.
-Recent efforts in open-vocabulary SGG (OVSGG) leverage Vision-Language Models (VLMs) [31, 32, 33] or MLLMs [8, 9, 34] to handle novel entities/relations, yet limited modeling of complex multi-entity interactions, and suffer from overfitting in base categories during SFT.
-Our Relation-R1 addresses OVSGG challenges by integrating a cognitive CoT-guided RL framework to model complex multi-entity interactions and mitigate SFT overfitting, enabling robust zero-shot reasoning without predefined category constraints.
+SGG is a critical task in scene understanding, with prior approaches falling into two groups: 1) Two-stage SGG, which sequentially detects objects and infers pairwise relations but suffers from error propagation (tang2020unbiased; zellers2018neural; li2022devil; li2023compositional; shi2025easy); 2) One-stage SGG, which unifies detection and relation prediction in end-to-end frameworks (e.g., DETR-based methods (carion2020end; li2022sgtr; chen2020simple)) but lacks multi-role involved relation modeling.
+Recent efforts in open-vocabulary SGG (OVSGG) employ Vision-Language Models (VLMs) (chen2024scene; li2023zero; li2025inter; chen2024expanding) or MLLMs (wang2024all; llava_spacesgg2025; li2024pixels) to handle novel entities/relations, yet limited modeling of complex multi-entity interactions, and suffer from overfitting in base categories during SFT.
+Our Relation-R1 addresses OVSGG challenges by adopting a cognitive CoT-guided RL framework to model complex multi-entity interactions and mitigate SFT overfitting, enabling robust zero-shot reasoning without predefined category constraints.
 Grounded Situation Recognition (GSR).
 GSR extends scene understanding by jointly modeling actions (verbs) and their associated semantic roles.
-Pratt et al.
-[10] pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization.
-Subsequent works [11, 35, 36, 37] improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations.
-Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution [24, 25, 26], potentially limiting real-world applicability.
-While LEX [38] realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions.
+Pratt et al(pratt2020grounded) pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization.
+Subsequent works (cho2021grounded; wei2021rethinking; cho2022collaborative; cheng2022gsrformer) improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations.
+Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution (tang2020unbiased; li2024nicest; shao2024knowledge; chen2023addressing), potentially limiting real-world applicability.
+While LEX (lei2024seeing) realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions.
 This paper proposes Relation-R1, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
-3 Methodology
-Figure 2: Overview of the proposed Relation-R1 framework.
+Figure 2: Relation-R1 framework.
 (a) SFT with progressive guidance of cognitive CoT, which enforces task-specific reasoning processes and format alignment.
-(b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-nary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
-3.1 Preliminaries
-Binary Relation Detection.
-This task requires the model to generate an open-ended scene graph caption given an image, explicitly marking semantic relationships between pairwise objects [8].
-The output is a triplet-contained caption annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates in < box> tags to localize the entity in the image, e.g., <ref>person</ref><box>[[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]]</box>.
+(b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-ary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
+Methodology
+Preliminaries.
+1) Binary Relation Detection: This task requires the model to generate an open-ended scene graph caption given an image, explicitly marking semantic relationships between pairwise objects (wang2024all).
+The output is a triplet-contained caption annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates (top-left/bottom-right) in < box>, e.g., <ref>person</ref><box>$[$$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]]</box>.
 Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes.
-N-ary Relation Detection.
-Similarly, this task aims to output a open-ended grounded situation frame [10] structured as a primary activity (i.e., multi-entity involved relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates.
+2) N-ary Relation Detection: It aims to output a open-ended grounded situation frame (pratt2020grounded) structured as a primary activity (i.e., multi-entity involved relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates.
 Each role explicitly defines the entity’s participation in the activity, e.g., <agent>person</agent><box>[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]</box>.
-3.2 Supervised Fine-Tuning
-As discussed in §1, directly applying GRPO encounters challenges in generating consistently formatted scene graph captions and grounded situation frame outputs.
+Supervised Fine-Tuning
+As discussed above, directly applying GRPO encounters challenges in generating consistently formatted scene graph captions and grounded situation frame outputs.
 To address this, we employ SFT as a foundational stage to establish task-specific output format adherence (c.f., Figure 2(a)).
-However, relying solely on the target answer format during SFT risks compromising the model’s underlying reasoning capabilities by encouraging a superficial alignment with formulated structures [13].
-To mitigate this potential issue, we introduce the cognitive chain-of-thought, a methodology that integrates stepwise cognitive processes such as object classification, object grounding, and visual relationship inference.
+However, relying solely on the target answer format during SFT risks compromising the model’s underlying reasoning capabilities by encouraging a superficial alignment with formulated structures (chu2025sft).
+To mitigate this potential issue, we introduce the cognitive chain-of-thought that integrates stepwise cognitive processes such as object classification, object grounding, and visual relationship inference.
 These cognitive CoTs are explicitly encapsulated within <think> tags, during SFT.
@@
 MLLM-generated CoT.
-For the general reasoning process, we design a cognitive CoT generation prompt that explicitly makes an extra strong MLLM (e.g., Qwen 2.5-VL [39]) to generate step-by-step reasoning processes.
-To be specific, this cognitive CoT generation prompt consists of a task definition, a ground-truth scene graph of the current image, and a CoT generation instruction.
+For the general reasoning process, we design a cognitive CoT generation prompt that explicitly makes an extra strong MLLM (e.g., Qwen 2.5-VL (Qwen2.5-VL)) to generate step-by-step reasoning processes.
+To be specific, this cognitive CoT generation prompt consists of a task definition, a ground-truth (GT) scene graph of the current image, and a CoT generation instruction.
 The former one makes the MLLM generate the correct task format.
 Meanwhile, the latter two anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors.
-The detailed cognitive CoT generation prompt is left in the appendix §B.
+The detailed cognitive CoT generation prompt is left in the Appendix.
 Progressive CoT Guidance.
-With a progressive paradigm that involves initial training on easy-to-learn fully template-based CoTs and subsequent fine-tuning on a small number of MLLM-generated CoTs, the MLLM gains the foundational capability to generate format-compliant outputs and grounded reasoning processes, which is critical for GRPO training.
-3.3 Group Relative Policy Optimization
-Following the Deepseek-R1 [12], the group relative policy optimization is adopted as the RL algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b).
-Distinct from critic-dependent methods [40], GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity [12].
+With a progressive paradigm that involves initial SFT training on easy-to-learn template-based CoTs and subsequent fine-tuning on a small number of MLLM-generated CoTs, the MLLM gains foundational capability to generate format-compliant outputs and grounded reasoning processes, contributing to GRPO training.
+Group Relative Policy Optimization
+Following the Deepseek-R1 (DeepseekR1), the group relative policy optimization is adopted as the RL algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b).
+Distinct from critic-dependent methods (schulman2017proximal), GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity (DeepseekR1).
 Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{1},o_{2},\ldots,o_{G}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge.
 The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging the following objectives:
-$J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$ (1)
+$J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}$ (1)
+$[\frac{1}{G}\!\sum_{i=1}^{G}\!\min\left(\!\rho_{i}A_{i},\text{clip}(\rho_{i},1\!-\!\epsilon,1\!+\!\epsilon)A_{i}\!\right)\!-\!\beta\!D_{\text{KL}}(\!\pi_{\theta}\!\|\!\pi_{\text{ref}})],$
 where $\rho_{i}$ = $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ quantifies policy change, $\epsilon$ controls clipping thresholds, and $\beta$ penalizes deviations via the KL divergence term.
 The advantage score $A_{i}$ is standardized as:
-$\small A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
-here $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
-$\small D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1.$ (3)
+$A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
+here $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy:
+$D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1.$ (3)
 This term ensures controlled exploration without excessive divergence from $\pi_{\theta_{\text{ref}}}$.
 The reward $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding.
-It incorporates rule-based rewards, which consist of three crucial components: Format compliance rewards, binary relation rewards, and N-nary relation rewards.
+It incorporates rule-based rewards, which consist of three crucial components: Format compliance rewards, binary relation rewards, and N-ary relation rewards.
 The former enforces syntactic validity by penalizing outputs that deviate from predefined structural constraints.
@@
 Format Rewards. This reward enforces strict adherence to the reasoning template format: <think> thinking process </think><answer> answer </answer>. The reward function is defined as:
-$\small r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
+$r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
 This binary mechanism ensures syntactic compliance while allowing content flexibility.
 Binary Relation Rewards.
-Inspired by established evaluation practices [17, 8], we devise the sample-level triplet recall $R$ and sample-level mean recall $mR$ in the response of each sample.
+Inspired by established evaluation practices (yang2022panoptic; wang2024all), we devise the sample-level triplet recall $R$ and sample-level mean recall $mR$ in the response of each sample.
 The former is the ratio of correctly predicted triplets to the total ground-truth triplets in a sample; the latter is the average recall across all predicate categories.
@@
 The final binary relation reward is computed as a weighted combination of these two metrics:
-$\small r_{\text{binary}}(o)=\alpha\cdot R+(1-\alpha)\cdot mR,$ (5)
+$r_{\text{binary}}(o)=\alpha\cdot R+(1-\alpha)\cdot mR,$ (5)
 where $\alpha$ balances the trade-off between two metrics.
 N-ary Relation Rewards.
-Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary relation rewards, which assess the model’s ability to detect primary activity (verb) and multi-entity interactions.
-Specifically, we calculate the entity value $V_{e}$ and grounded value $V_{grnd}$ for each sample [10].
-These respectively represent the correct proportion in predicted entity categories and roles, and the spatial localization of those entities:
-  • Entity Accuracy: An entity within an event is considered correct if its predicted category and semantic role align with the ground truth, and the predicted activity matches the ground-truth.
-  • Spatial Localization: The predicted bounding box of each role obtains IoU >=0.5 with the ground-truth’s bounding box.
+Similarly, to test higher-order relational structures in grounded situation frames, we devise N-ary relation rewards to assess the model’s ability to detect primary activity (verb) and multi-entity interactions.
+Specifically, we calculate the entity value $V_{e}$ and grounded value $V_{grnd}$ for each sample (pratt2020grounded).
+They represent the correct proportion in predicted entity classes and roles, and the spatial localization of those entities, respectively:
+  • Entity Accuracy: An entity within an event is considered correct if its predicted category and semantic role align with the GT, and the predicted activity matches the GT.
+  • Spatial Localization: The predicted bounding box of each role obtains IoU >=0.5 with the GT’s bounding box.
 The eventual N-ary relation reward is computed as a weighted combination of these two accuracies:
-$\small r_{\text{n-ary}}(o)=\beta\cdot V+(1-\beta)\cdot V_{grnd},$ (6)
+$r_{\text{n-ary}}(o)=\beta\cdot V_{e}+(1-\beta)\cdot V_{grnd},$ (6)
 where the hyperparameter $\beta$ balances the trade-off between the two values.
@@
 During training, we design a multi-task gate that dynamically selects task-specific rewards based on the presence of <ref> tags in the solutions.
-4 Experiment
-4.1 Experimental Setup
-Datasets.
-1) Panoptic Scene Graph (PSG): The official PSG dataset [17], built upon the COCO dataset [41], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing.
-It integrates 80 “thing” object categories and 53 “stuff” object categories, alongside 56 relation categories.
-Notably, ASMv2 [8] reconstructs the dataset into a scene graph caption format, utilizing a split of 42,250 training images and 1,000 evaluation images.
-To mark fair comparison, we followed ASMv2 [8], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training.
-Besides, we also present results using the conventional scene graph format, i.e., a list of visual triplets structured as [subject, subject bounding box, object, object bounding box, relation] on the official dataset [^1] More details of the scene graph format are in the appendix §C..
-2) SWiG: The SWiG dataset [10] extends the imSitu dataset [42] by adding extra bounding box annotations for semantic roles, with 69.3% of semantic roles localized.
-Each image is annotated with a verb and a variable number of semantic roles (1$\sim$6 roles per image), where each verb is associated with three verb frames annotated by independent human annotators.
-The dataset comprises 25,200 testing images spanning 504 verb categories, 190 semantic role categories, and 9,929 noun entity categories.
-Unlike traditional output structures that merely predict categories and bounding boxes, we constructed grounded situation frames by filling verb-centric templates (e.g., “The AGENT drinks a LIQUID from a CONTAINER”) with three critical components: role tags (e.g., <agent>, entity categories (e.g., child), and coordinates of corresponding bounding boxes in <box> tags.
+Experiment
+[…]
+  Method Size Recall mRecall Mean
+Close-ended
+IMP (xu2017scene) - 16.50 6.50 11.50
+MOTIFS (zellers2018neural) - 20.00 9.10 14.55
+VCTree (tang2019learning) - 20.60 9.70 15.15
+GPSNet (lin2020gps) - 17.80 7.00 12.4
+PSGFormer (yang2022panoptic) - 18.60 16.70 17.65
+Open-ended
+TextPSG (Zhao_2023_ICCV) - 4.80 - -
+R1-SGG (chen2025compile) 2B 27.83 17.03 22.43
+R1-SGG (chen2025compile) 7B 28.77 17.55 23.16
+Relation-R1(Ours) 3B 25.87 21.32 23.60
+ASMv2 (wang2024all) 13B 14.20 10.30 12.23
+SpaceSGG (llava_spacesgg2025) 13B 15.43 13.23 14.33
+Relation-R1 (Ours) 3B 22.33 20.07 21.20
+[…]
 Evaluation Metrics.
-For binary relation detection, we followed the standard SGG evaluation protocol [24, 8] 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object.
-2) mean Recall (mRecall): The average recall across all predicate categories.
-As for N-ary relation detection, we adopted the same five evaluation metrics as [10]: 1) verb: the accuracy of verb prediction.
+For binary relation detection, we followed the standard SGG evaluation protocol (tang2020unbiased; wang2024all) 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, with both label accuracy and bounding box IoU >= 0.5 for subject and object.
+2) mean Recall (mRecall): The average recall across all predicate classes.
+For N-ary relation detection, we adopted the same evaluation metrics as (pratt2020grounded): 1) verb: the accuracy of verb prediction.
 2) value: the noun accuracy for individual semantic role.
@@
 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
-Implementation Details. Due to space constraints, details are provided in the appendix §A.
-[…]
-         Method Model Size Format Recall mRecall Mean
-Close-ended
-IMP [16] ${}_{\text{CVPR'17}}$ 16.50 6.50 11.50
-MOTIFS [25] ${}_{\text{CVPR'18}}$ 20.00 9.10 14.55
-VCTree [43] ${}_{\text{CVPR'19}}$ 20.60 9.70 15.15
-GPSNet [44] ${}_{\text{CVPR'20}}$ 17.80 7.00 12.4
-PSGFormer [17] ${}_{\text{ECCV'22}}$ 18.60 16.70 17.65
-Open-ended
-TextPSG [45] ${}_{\text{ICCV'23}}$ Standard Scene Graph 4.80 - -
-R1-SGG [46] ${}_{\text{ArXiv'25}}$ 2B 27.83 17.03 22.43
-R1-SGG [46] ${}_{\text{ArXiv'25}}$ 7B 28.77 17.55 23.16
-Relation-R1 (Ours) 3B 25.87 21.32 23.60
-ASMv2 [8] ${}_{\text{ECCV'24}}$ 13B Scene Graph Caption 14.20 10.30 12.23
-LLaVA-SpaceSGG [9] ${}_{\text{WACV'25}}$ 13B 15.43 13.23 14.33
-Relation-R1 (Ours) 3B 22.33 20.07 21.20
-Table 2: Performance(%) comparison with various GSR models on the SWiG [10] dataset.
+Implementation Details. Refer to the Appendix.
          Method Answer Verb Value Value-all Grnd Grnd-all
-ISL [10] ${}_{\text{ECCV'20}}$ Close-ended 39.36 30.09 18.62 22.73 7.72
-JSL [10] ${}_{\text{ECCV'20}}$ 39.94 31.44 18.87 24.86 9.66
-GSRTR [11] ${}_{\text{BMVC'21}}$ 40.63 32.15 19.28 25.49 10.10
-CoFormer [36] ${}_{\text{CVPR'22}}$ 44.66 35.98 22.22 29.05 12.21
-SituFormer [35] ${}_{\text{AAAI'22}}$ 44.20 35.24 21.86 29.22 13.41
-GSRFormer [37] ACM ${}_{\text{ACM MM'22}}$ 46.53 37 .48 23.32 31.53 14.23
-OpenSU [47] ${}_{\text{ICCVW'23}}$ Open-ended 50.10 41.20 26.56 34.27 15.70
+ISL (pratt2020grounded) $_{\text{ECCV'20}}$ Close-ended 39.36 30.09 18.62 22.73 7.72
+JSL (pratt2020grounded) $_{\text{ECCV'20}}$ 39.94 31.44 18.87 24.86 9.66
+GSRTR (cho2021grounded) $_{\text{BMVC'21}}$ 40.63 32.15 19.28 25.49 10.10
+CoFormer (cho2022collaborative) $_{\text{CVPR'22}}$ 44.66 35.98 22.22 29.05 12.21
+SituFormer (wei2021rethinking) $_{\text{AAAI'22}}$ 44.20 35.24 21.86 29.22 13.41
+GSRFormer (cheng2022gsrformer) $_{\text{ACM MM'22}}$ 46.53 37 .48 23.32 31.53 14.23
+OpenSU (liu2023opensu) $_{\text{ICCVW'23}}$ Open-ended 50.10 41.20 26.56 34.27 15.70
 Relation-R1 (Ours) 57.26 46.66 30.92 40.21 30.18
-4.2 Comparison with State-of-the-Art Methods
-We compared our Relation-R1 for unified relation comprehension with the existing state-of-the-art binary relation detection and N-ary relation detection approaches in Table 1 and Table 2, respectively.
-[…]
-[…]
-4.3 Diagnostic Experiment
-To gain more insights and reduce the mutual influence between the two tasks during training, we conducted a set of comprehensive studies on the Relation-R1 trained with binary and N-ary relation detection, respectively.
-Table 3: Performance comparison (%) across various cognitive CoT strategies trained on single binary relation detection and $N$-ary relation detection, separately.
-Blue indicates metrics without correct verb constraints.
-  Binary Relation $N$-ary Relation
+Table 2: Performance(%) comparison with various GSR models on the SWiG (pratt2020grounded) dataset.
+  Binary Relation NN-ary Relation
 Method CoT Recall mRecall Verb Value Value-all Grnd Grnd-all
@@
 SFT + RL Progressive 22.57 20.57 71.04 61.26 (78.19) 44.98 (49.73) 49.43 (61.25) 36.09 (42.35)
+Table 3: Performance comparison (%) across various cognitive CoT strategies trained on single binary relation detection and NN-ary relation detection, separately.
+Blue indicates metrics without correct verb constraints.
+Figure 3: The statistics of binary and NN-ary relation rewards.
+Comparison with State-of-the-Art Methods
+We compared Relation-R1 for unified relation comprehension with the existing state-of-the-art binary relation detection (Table 1) and N-ary relation detection approaches in (Table 2).
+[…]
+Figure 4: Qualitative results of binary relation detection.
+[…]
+Diagnostic Experiment
+To gain more insights and reduce the mutual influence between the two tasks during training, we conducted a set of comprehensive studies on the Relation-R1 trained with binary and N-ary relation detection, respectively.
 Cognitive Chain-of-Thoughts.
@@
 These results highlight the superior ability of the progressive CoT to guide complex relation detection.
-Figure 3: The statistics of binary relation rewards.
-Figure 4: The statistics of $N$-ary relation rewards.
 Reward Analysis.
-We visualized the reward variation curves for binary and N-ary relation detection in Figure 3 and Figure 4, respectively.
-It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT.
+We visualized the reward variation curves for binary and N-ary relation detection in Figure 3, respectively.
+It can be seen that at the beginning stage, both tasks obtains decent rewards due to the foundation laid by SFT.
 With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend.
 In binary relation detection, guided with progressive CoT yields a slightly enhanced overall reward trajectory, although the improvement is not significant.
-This might be since the gap between the CoT generated by MLLM (e.g., Qwen [39]) and that produced by the template is not obvious.
-Conversely, N-ary relation detection exhibits a significant surge in reward, primarily attributed to the model’s emergent capability to generate synonymous grounded situation frames (c.f.
-§5).
+This might be since the gap between the CoT generated by MLLM (e.g., Qwen (Qwen2.5-VL)) and that produced by the template is not obvious.
+Conversely, N-ary relation detection exhibits a significant surge in reward, primarily attributed to the model’s emergent capability to generate synonymous grounded situation frames.
 Completion Length Analysis.
-Figure 5: The statistics of binary relation completion lengths.
-Figure 6: The statistics of $N$-ary relation completion lengths.
-The evolution of completion length during GRPO training is depicted in Figure 5 and Figure 6.
+Figure 5: Completion lengths of binary and NN-ary relation.
+The statistic of completion length during GRPO training is in Figure 5.
 Consistent with the reward trends, the completion lengths for binary relation detection exhibit a gradual increase as GRPO progresses, suggesting the generation of more detailed and descriptive outputs over time.
@@
 The progressive CoT guidance, by enabling the model to explore synonymous relational expressions and incorporate more nuanced details, leads to an increase in the average completion length.
-Table 4: Few-Shot GSR performance comparison.
-         Method Setting Verb Value Value-all Grnd Grnd-all
-GSRTR [11] ${}_{\text{BMVC'21}}$ One-Shot 0.35 0.10 0.01 0.08 0.00
-CoFormer [36] ${}_{\text{CVPR'22}}$ 0.63 0.23 0.02 0.14 0.01
-Relation-R1 (Ours) 31.67 15.37 4.15 14.01 3.41
-$w$/$o$ RL 27.31 10.21 1.43 9.35 1.22
-GSRTR [11] ${}_{\text{BMVC'21}}$ 4-Shot 6.95 4.32 1.83 2.79 0.65
-CoFormer [36] ${}_{\text{CVPR'22}}$ 12.39 8.27 3.93 5.28 1.25
-Relation-R1 (Ours) 38.02 21.87 7.39 19.64 5.92
-$w$/$o$ RL 34.88 17.14 3.94 15.63 3.38
-Few-Shot N-ary Relation Detection.
-We also evaluated Relation-R1’s few-shot generalization capability (Table 4).
-In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Grnd-all) compared to CoFormer’s [36] 0.63% and 0.01%, respectively.
-This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Grnd-all), significantly surpassing previous methods.
-These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
-Joint Binary and N-ary Relation Detection.
-Table 5: Performance comparison (%) across various cognitive CoT strategies trained on joint binary relation detection and $N$-ary relation detection, separately.
-Blue indicates metrics without correct verb constraints.
-  Binary Relation $N$-ary Relation
-Method CoT Recall mRecall Verb Value Value-all Grnd Grnd-all
-SFT - 14.89 12.66 56.23 45.81 (66.00) 30.31 (34.06) 38.59 (54.11) 28.77 (36.18)
-SFT + RL Template-based 22.33 20.07 57.26 46.66 (66.14) 30.92 (34.75) 40.21 (55.50) 30.18 (37.64)
-SFT + RL MLLM-generated 19.59 18.30 53.00 42.27 (65.11) 26.67 (34.65) 35.69 (55.36) 25.89 (37.61)
-SFT + RL Progressive 27.19 26.12 73.25 66.18 (82.54) 52.90 (57.83) 55.20 (67.13) 42.36 (48.81)
-Beyond training Relation-R1 solely on binary or N-ary relation detection tasks (as discussed in §3), we also investigated the joint training of both tasks within the same Relation-R1 model.
-The results of this joint training are presented in Table 5333The Grnd-all metric excludes ground-truth bounding boxes represented as [-1, -1, -1, -1] for calculation..
-From the analysis of these results, three primary conclusions can be drawn: 1) When utilizing both the SFT and RL strategies, our Relation-R1 consistently achieves higher performance across nearly all metrics for both binary and N-ary relation detection tasks.
-Similar to single-task training, the MLLM-generated CoT strategy underperforms in the N-ary task.
-This is likely due to the inherent generality of the CoT generated by Qwen [39], which may not precisely align with the ground-truth annotations, consequently leading to lower Verb accuracy.
-2) Our proposed progressive CoT-guided paradigm consistently achieves the highest performance across all evaluated metrics for both tasks in the joint learning setting.
-3) The joint progressive CoT-guided training across both binary and N-ary tasks fosters a synergistic mutual promotion effect.
-This is evidenced by the notable gains observed: for instance, a Recall of 27.19% in Table 5 vs.
-22.57% in Table 3 (for binary relation detection), and a Verb accuracy of 73.25% in Table 5 vs.
-71.04% in Table 3 (for N-ary relation detection).
-5 Qualitative Results
-Figure 7: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG [17] dataset. The orange area denotes the predicate distribution of the training set.
-Figure 8: Qualitative results of binary relation detection.
-Figure 9: Qualitative results of N-ary relation detection.
-Figure 10: Qualitative results of synonymous in N-ary relation detection.
-Comparison over All Predicates.
-We visualized predicate-level comparison against ASMv2 [8] and the predicate distribution in Figure 7.
-As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates.
-These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
-Binary and N-ary Relation Detection Results.
-We visualized some examples of binary and N-ary relation detection in Figure 8 and Figure 9, respectively.
-As seen, for binary relation detection (c.f.
-Figure 8), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference.
-For instance, Relation-R1 can first focus on the main objects (e.g., “cat” and “train”) and then infer their positions and relationships.
-Similarly, for N-ary relation detection (c.f.
-Figure 9), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box.
-This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
-Synonymous Visualization.
-Figure 10 displays synonymous results observed in N-ary relation detection.
-As evinced by these results, our progressive CoT-guided training paradigm empowers the model to generate diverse synonymous expressions, encompassing entities, semantic roles, and activities.
-For instance, in Figure 10(top), Relation-R1 demonstrates the ability to output synonymous agents such as “person” and “man”.
-Furthermore, the model can generate semantically similar and plausible activities, exemplified by “barbecues” and “grills” (c.f.
-Figure 10(bottom)).
-Even within the same activity “rests”, it can differentiate and utilize varying semantic role names like “item” and “bodypart” (c.f.
-Figure 10(row 2)).
-These synonymous outputs indicate a more flexible and robust semantic understanding beyond rigid lexical matching, thereby significantly enhancing the model’s generalization capabilities to diverse linguistic formulations and real-world variations.
-6 Conclusion
-This paper reveals critical limitations in MLLM visual relation understanding, particularly for N-ary relationships and language-induced bias.
-We propose Relation-R1, a unified framework with a two-stage approach: establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, followed by RL with GRPO for refinement, promoting visual-semantic grounding and robust generalization.
-Furthermore, progressively applied cognitive CoT guidance, transitioning from template-based to MLLM-generated reasoning, further enhances generalization, especially for synonymous N-ary relations.
-Extensive experiments on PSG and SWiG validate Relation-R1’s effectiveness, demonstrating SOTA performance in both binary and N-ary relation detection.
-Integrating progressive CoT-guided SFT with R1-inspired RL refinement, Relation-R1 advances MLLMs towards more human-like relational reasoning.
+Qualitative Results.
+Figure 4 shows a qualitative example of binary relation detection.
+As seen, the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference.
+For instance, after identifying the central bed, Relation-R1 pinpoints related objects e.g., rug and chair, and explicitly grounds the inferred spatial relations (on, beside) between them.
+Our approach demonstrates the successful decomposition of the scene into accurate, grounded object-relation pairs.
+Conclusion
+This paper reveals key limitations in MLLM visual relation understanding, particularly for N-ary relations and language-induced bias.
+We propose Relation-R1, a unified two-stage framework: build a reasoning foundation and structured outputs via cognitive CoT-guided SFT, followed by GRPO, promoting visual-semantic grounding and robust generalization.
+Besides, progressively applied cognitive CoT guidance, transitioning from template-based to MLLM-generated reasoning, further enhances generalization, especially for synonymous N-ary relations.
+Experiments on PSG and SWiG validate Relation-R1’s effectiveness in both binary and N-ary relation detection.
+Employing progressive CoT-guided SFT with R1-inspired refinement, Relation-R1 advances MLLMs towards more human-like relational reasoning.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China Young Scholar Fund (62402408), the Hong Kong SAR RGC General Research Fund under Grant (16208823), and the Hong Kong SAR RGC Early Career Scheme (26208924).
+This research was partially conducted by ACCESS – AI Chip Center for Emerging Smart Systems, supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government.
+This work was also supported by the Hong Kong SAR RGC General Research Fund under Grant (16208823) and the Hong Kong SAR RGC Early Career Scheme (26208924).
+Summary of the Appendix
+To enhance the understanding of the main paper, supplementary material containing further details are provided, organized as follows:
+  • §A elaborates on the implementation details of Relation-R1.
+  • §B introduces the cognitive CoT generation prompt.
+  • §C details the scene graph format utilized in this work.
+  • §D provides additional experimental results.
+  • §E shows further qualitative results.
+  • §F discusses limitations and broader impacts of Relation-R1.
 Appendix A Implementation Details
-We utilized Qwen2.5-VL-3B [39] as our base model.
+We utilized Qwen2.5-VL-3B (Qwen2.5-VL) as our base model.
 Relation-R1 was trained on an 8xA100 80GB GPU server.
@@
 The learning rates were set to 2e-5 for SFT and 1e-6 for GRPO, respectively.
-Qwen2.5-VL-72B [39] served as the CoT generator for producing MLLM-generated CoTs.
+Qwen2.5-VL-72B (Qwen2.5-VL) served as the CoT generator for producing MLLM-generated CoTs.
+The weighting hyperparameters $\alpha$ and $\beta$ were all set to 0.5 in the reward function.
 Progressive training began with supervised fine-tuning on binary and N-ary relation detection tasks for 2 epochs using template-based CoT, followed by fine-tuning on 4k MLLM-generated CoT samples.
@@
 Figure S2: Grounded situation frame CoT generation prompt.
-As mentioned in §3.2, we employ two distinct prompts (Figure S1 and Figure S2) to guide cognitive CoT generation for binary and N-ary relation detection, respectively.
+As mentioned above, we employ two distinct prompts (Figure S1 and Figure S2) to guide cognitive CoT generation for binary and N-ary relation detection, respectively.
 Each prompt strategically comprises: 1) a concise task definition outlining the detection objective; 2) an illustrative in-context example demonstrating the desired reasoning steps and output format; and 3) a specific instruction prompting the model to generate CoT of the current image.
@@
 The scene graph format prompt employed for GRPO training is displayed in Figure S3.
-Furthermore, to enable the model’s generation of think process, we utilize templated-CoT (c.f.
-§3.2) for supervised fine-tuning.
-Appendix D Discussions
+Furthermore, to enable the model’s generation of the think process, we utilize template-based CoT for supervised fine-tuning.
+Appendix D More Experimental Results
+Table S1: Few-Shot GSR performance comparison.
+         Method Setting Verb Value Value-all Grnd Grnd-all
+GSRTR (cho2021grounded) $_{\text{BMVC'21}}$ One-Shot 0.35 0.10 0.01 0.08 0.00
+CoFormer (cho2022collaborative) $_{\text{CVPR'22}}$ 0.63 0.23 0.02 0.14 0.01
+Relation-R1 (Ours) 31.67 15.37 4.15 14.01 3.41
+ww/$o$ RL 27.31 10.21 1.43 9.35 1.22
+GSRTR (cho2021grounded) $_{\text{BMVC'21}}$ 4-Shot 6.95 4.32 1.83 2.79 0.65
+CoFormer (cho2022collaborative) $_{\text{CVPR'22}}$ 12.39 8.27 3.93 5.28 1.25
+Relation-R1 (Ours) 38.02 21.87 7.39 19.64 5.92
+ww/$o$ RL 34.88 17.14 3.94 15.63 3.38
+Few-Shot N-ary Relation Detection.
+We also evaluated Relation-R1’s few-shot generalization capability (Table S1).
+In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Grnd-all) compared to CoFormer’s (cho2022collaborative) 0.63% and 0.01%, respectively.
+This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Grnd-all), significantly surpassing previous methods.
+These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
+Joint Binary & N-ary Relation Detection.
+Table S2: Performance comparison (%) across various cognitive CoT strategies trained on joint binary relation detection and NN-ary relation detection, separately.
+Blue indicates metrics without correct verb constraints.
+  Binary Relation NN-ary Relation
+Method CoT Recall mRecall Verb Value Value-all Grnd Grnd-all
+SFT - 14.89 12.66 56.23 45.81 (66.00) 30.31 (34.06) 38.59 (54.11) 28.77 (36.18)
+SFT + RL Template-based 22.33 20.07 57.26 46.66 (66.14) 30.92 (34.75) 40.21 (55.50) 30.18 (37.64)
+SFT + RL MLLM-generated 19.59 18.30 53.00 42.27 (65.11) 26.67 (34.65) 35.69 (55.36) 25.89 (37.61)
+SFT + RL Progressive 27.19 26.12 73.25 66.18 (82.54) 52.90 (57.83) 55.20 (67.13) 42.36 (48.81)
+Beyond solely training Relation-R1 on binary or N-ary relation detection tasks, we also investigated the joint training of both tasks within the same Relation-R1 model.
+The results of this joint training are presented in Table S2333The Grnd-all metric excludes ground-truth bounding boxes represented as [-1, -1, -1, -1] for calculation..
+From the analysis of these results, three primary conclusions can be drawn: 1) When utilizing both the SFT and RL strategies, our Relation-R1 consistently achieves higher performance across nearly all metrics for both binary and N-ary relation detection tasks.
+Similar to single-task training, the MLLM-generated CoT strategy underperforms in the N-ary task.
+This is likely due to the inherent generality of the CoT generated by Qwen (Qwen2.5-VL), which may not precisely align with the ground-truth annotations, consequently leading to lower Verb accuracy.
+2) Our proposed progressive CoT-guided paradigm consistently achieves the highest performance across all evaluated metrics for both tasks in the joint learning setting.
+3) The joint progressive CoT-guided training across both binary and N-ary tasks fosters a synergistic mutual promotion effect.
+This is evidenced by the notable gains observed: for instance, a Recall of 27.19% in Table S2 vs.
+22.57% in Table 3 (for binary relation detection), and a Verb accuracy of 73.25% in Table S2 vs.
+71.04% in Table 3 (for N-ary relation detection).
+Appendix E Qualitative Results
+Figure S4: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG (yang2022panoptic) dataset.
+The orange area denotes the predicate distribution of the training set.
+Figure S5: Qualitative results of binary relation detection.
+Figure S6: Qualitative results of N-ary relation detection.
+Figure S7: Qualitative results of synonymous in N-ary relation detection.
+Comparison over All Predicates.
+We visualized predicate-level comparison against ASMv2 (wang2024all) and the predicate distribution in Figure S4.
+As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates.
+These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
+Binary and N-ary Relation Detection Results.
+We visualized some examples of binary and N-ary relation detection in Figure S5 and Figure S6, respectively.
+As seen, for binary relation detection (c.f.
+Figure S5), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference.
+For instance, Relation-R1 can first focus on the main objects (e.g., “cat” and “train”) and then infer their positions and relationships.
+Similarly, for N-ary relation detection (c.f.
+Figure S6), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box.
+This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
+Synonymous Visualization.
+Figure S7 displays synonymous results observed in N-ary relation detection.
+As evinced by these results, our progressive CoT-guided training paradigm empowers the model to generate diverse synonymous expressions, encompassing entities, semantic roles, and activities.
+For instance, in Figure S7(top), Relation-R1 demonstrates the ability to output synonymous agents such as “person” and “man”.
+Furthermore, the model can generate semantically similar and plausible activities, exemplified by “barbecues” and “grills” (c.f.
+Figure S7(bottom)).
+Even within the same activity “rests”, it can differentiate and utilize varying semantic role names like “item” and “bodypart” (c.f.
+Figure S7(row 2)).
+These synonymous outputs indicate a more flexible and robust semantic understanding beyond rigid lexical matching, thereby significantly enhancing the model’s generalization capabilities to diverse linguistic formulations and real-world variations.
+Appendix F Discussions
 Limitations.